### v1

# Improved polynomial rates of memory loss for nonstationary intermittent dynamical systems

Abstract.
We study nonstationary dynamical systems formed by sequential concatenation of nonuniformly expanding maps with a uniformly expanding first return map. Assuming a polynomially decaying upper bound on the tails of first return times that is nonuniform with respect to location in the sequence, we derive a corresponding sharp polynomial rate of memory loss. As applications, we obtain new estimates on the rate of memory loss for random ergodic compositions of Pomeau–Manneville type intermittent maps and intermittent maps with unbounded derivatives.
Date: October 28, 2024
A.K. is supported by EPSRC grant EP/V053493/1. J.L. is supported by JSPS via the project LEADER. The authors thank Wael Bahsoun for useful discussions, as well as Loughborough University and Tokai University for hospitality over the course of this work.
1. Introduction
In this paper, we study statistical properties of nonstationary intermittent dynamical systems. By nonstationarity we mean that the rules governing the system’s evolution are not fixed, but they vary with time under the influence of noise, fluctuating environment, control signals, or other external factors. This is modeled by time-dependent compositions $T_{n}\circ\cdots\circ T_{1}$ along a sequence of transformations $T_{i}:X\to X$ acting on a state space $X$, where each $T_{i}$ describes the rules governing the evolution of states at time $i$. When an initial state $x_{0}\in X$ is sampled from a random distribution, the successive states $x_{n}:=T_{n}\circ\cdots\circ T_{1}(x)$ form a random process that is generally nonstationary. This is in contrast with the conventional setting of ergodic theory developed in the stationary context of measure preserving transformations. The statistical theory of nonstationary dynamical systems characterizes the long term behavior of $x_{n}$, for example through limit theorems for partial sums $\sum_{k=0}^{n-1}v_{k}(x_{k})$ generated by observations $v_{k}:X\to{\mathbb{R}}$ along the time-dependent trajectory.
[…]
Intermittent dynamical systems are chaotic systems that include laminar regions where trajectories can become trapped for long periods of time before transitioning to more chaotic behavior, as seen in the Pomeau–Manneville scenario [24]. The rate of memory loss for such systems, whether stationary or nonstationary, is known to be polynomial [1, 29, 18, 16].
[…]
$|(T_{1,n})_{*}\mu-(T_{1,n})_{*}\nu|=O(n^{-\gamma})$ (1.1)
[…]
We apply our abstract result (Theorem 2.6) to obtain new bounds on memory loss for random ergodic compositions of intermittent interval maps. These results can be viewed as complementary to those of Bahsoun, Bose, and Ruziboev [2], who derived estimates on the quenched decay rate of correlations for random i.i.d. compositions of intermittent maps. Their approach, which differs from ours, was based on the construction of a random Young tower in the spirit of [3]. Besides the standard example due to Liverani, Saussol, and Vaienti [18], we present applications for two families of intermittent maps with unbounded derivatives (see Figure 1), studied by Cristadoro, Haydn, Marie, and Vaienti in [4], and more recently by Muhammad and Ruziboev in [20]: Pikovsky maps [23] and Grossmann–Horner maps [12]. In our analysis we exploit the distortion bounds proved in [20].
001/2121/21 / 21111001111
11-1- 100111111-1- 1001111
11-1- 100111111-1- 1001111
Figure 1. Graphs of three intermittent maps. Left: Cui’s map in (3.4); Middle: Pikovsky map (3.7); Right: Grossmann–Horner map (3.14).
In the remainder of this introduction, we illustrate our main theorem by applying it to the example from [18]. That is, for $\gamma\in(0,1)$, we consider the map $T:X\to X$ on the unit interval $X=[0,1]$ defined by
$T(x)=\begin{cases}x(1+2^{\gamma}x^{\gamma}),&x\in[0,\tfrac{1}{2}),\\ 2(x-\tfrac{1}{2}),&x\in[\tfrac{1}{2},1].\end{cases}$ (1.2)
Let $\beta\in(0,1)$ and $a_{\beta}>2^{\beta}(\beta+2)$. Following [18], we define the cone
${\mathcal{C}}_{*}(\beta)=\{f\in C((0,1])\cap L^{1}\,:\,$ $f\geq 0$, $f$ decreasing,
[…]
This cone contains densities of invariant absolutely continuous probability measures for maps (1.2) with parameters $\gamma\in[0,\beta]$.
Let $(T_{n})$ be a sequence of maps $T_{1},T_{2},\ldots$ in the family (1.2) corresponding to parameters $\gamma_{1},\gamma_{2},\ldots$, and suppose that $\sup_{n}\gamma_{n}\leq\gamma_{*}<1$. By [16, Theorem 1.1], whenever $\mu$ and $\nu$ are measures with Hölder continuous densities,
$|(T_{1,n})_{*}(\mu-\nu)|=O(n^{-\gamma_{*}}),\quad\text{as $n\to\infty$.}$ (1.3)
[…]
Theorem 1.1.
Let $\gamma\in(0,1)$. There exists a constant $\kappa_{0}\in(0,1)$ such that the following holds. Assume that for some $a>0$, $\kappa\leq\kappa_{0}$, and $N\in{\mathbb{N}}$,
#$\frac{\#\{1\leq k\leq n\>:\>\gamma_{k}\leq\gamma\}}{n}\in[a(1-\kappa),a(1+\kappa)]$ (1.4)
holds for all $n\geq N$. Let $\mu$ and $\mu^{\prime}$ be probability measures with Hölder continuous densities, and let $\nu$ be a probability measure on $X$ with density in ${\mathcal{C}}_{*}(\gamma)$. Then
  • (a)
    $|(T_{1,n})_{*}(\mu-\nu)|=O(n^{-1/\gamma+1})$,
  • (b)
    $|(T_{1,n})_{*}(\mu-\mu^{\prime})|=O(n^{-1/\gamma})$.
Remark 1.2.
A concrete, although complicated, expression for $\kappa_{0}$ depending on dynamical constants could be extracted from the proofs of Propositions 3.1 and 4.6.
Remark 1.3.
Condition (1.4) is satisfied by random ergodic compositions of maps in (1.2) provided that maps with parameters below $\gamma$ are sampled with positive probability. See Corollary 3.3 for a precise statement.
Structure of the paper
In Section 2, we present our main result, an abstract version of Theorem 1.1. In Section 3, we formulate and prove the applications for intermittent maps mentioned in the Introduction. Section 4 contains the proof of the main result, and Appendix A contains the proof of a proposition stated Section 2 that is used in one of the applications.
2. Abstract setup and results
2.1. Nonstationary nonuniformly expanding dynamical systems, nonuniform tail bounds
Let $(X,d)$ be a bounded metric space. We endow $X$ with the Borel sigma-algebra, and we only work with measurable sets. For each $k\geq 1$, let $Y_{k}\subset X$ and let $m_{k}$ be a probability measure on $X$ with $m_{k}(Y_{k})=1$. We consider a sequence $T_{1},T_{2},\ldots$ of transformations $T_{i}:X\to X$ and denote $T_{k,\ell}=T_{\ell}\circ\cdots\circ T_{k}$. (If $k>\ell$, then $T_{k,\ell}$ is the identity map.)
For a nonnegative measure $\mu$ on $Y_{k}$ with density $\rho=d\mu/dm_{k}$, we denote by $|\mu|_{{\mathrm{LL}},k}$ the Lipschitz seminorm of the logarithm of $\rho:Y_{k}\to{\mathbb{R}}_{+}$:
$|\mu|_{{\mathrm{LL}},k}=\sup_{y\neq y^{\prime}\in Y_{k}}\frac{|\log\rho(y)-\log\rho(y^{\prime})|}{d(y,y^{\prime})},$
with a convention that $\log 0=-\infty$ and $\log 0-\log 0=0$.
We suppose that there exist constants $\lambda>1$, $K>0$, $\delta_{0}>0$ and $n_{0}\geq 1$ such that the following assumptions hold.
For $x\in X$, $k\geq 1$, define return time functions $\tau_{k}:X\to\{1,2,\ldots\}$,
$\tau_{k}(x)=\inf\{n\geq 1:T_{k,k+n-1}(x)\in Y_{k+n}\}.$
First, for each $k\geq 1$, we assume existence of a finite or countable partition ${\mathcal{P}}_{k}$ of $X$, such that $Y_{k}$ is ${\mathcal{P}}_{k}$-measurable. Moreover, for each $a\in{\mathcal{P}}_{k}$ we make the following assumptions:
  1. (NU:1)
    $m_{k}(a)>0$ for all $a\subset Y_{k}$.
  2. (NU:2)
    $\tau_{k}$ is constant on $a$ with value $\tau_{k}(a)$.
  3. (NU:3)
    If $a\subset Y_{k}$, then the map $F_{a}=T_{k,k+\tau_{k}(a)-1}:a\to Y_{k+\tau_{k}(a)}$ is a bijection, and for all $y,y^{\prime}\in a$,
    $d(F_{a}(y),F_{a}(y^{\prime}))\geq\lambda d(y,y^{\prime}).$
    Further, $F_{a}$ is nonsingular with log-Lipschitz Jacobian with respect to $m_{k+\tau_{k}(a)}$:
    $\zeta=\frac{d(F_{a})_{*}(m_{k}|_{a})}{dm_{k+\tau_{k}(a)}}\quad\text{satisfies}\quad|\zeta|_{{\mathrm{LL}},k+\tau_{k}(a)}\leq K.$
  4. (NU:4)
    For all $x,x^{\prime}\in a$, with $F_{a}=T_{k,k+\tau_{k}(a)-1}$ as above,
    $\max_{0\leq j\leq\tau_{k}(a)}d(T_{k,k+j-1}(x),T_{k,k+j-1}(x^{\prime}))\leq Kd(F_{a}(x),F_{a}(x^{\prime})).$
In other words, the first return map $y\mapsto T_{k,k+\tau_{k}(y)-1}(y)$ with respect to the sequence $T_{k},T_{k+1},\ldots$ is full branch Gibbs-Markov, and returns from outside of $Y_{k}$ have bounded backward expansion.
To quantify mixing, for each $k\geq 1$ we define
$h^{k}(n)=m_{k}(\tau_{k}\geq n),$ (2.1)
and assume that there exist $\delta_{0}>1$ and $n_{0}\geq 0$ such that:
  1. (NU:5)
    $((T_{k,k+n-1})_{*}m_{k})(Y_{k+n})\geq\delta_{0}$ for every $k\geq 1$ and $n\geq n_{0}$.
Thus, our setting is a variant of that in [16], with the crucial difference that we relax the assumption on tail bounds in [16], allowing them to depend on $k$.
It is sometimes convenient to use the following to verify (NU:5):
Proposition 2.1.
Suppose that (NU:1-4) hold. Then, in order to satisfy (NU:5), it is sufficient that the following conditions hold:
  • (i)
    there exists $N_{\#}\geq 1$ such that $\sup_{k,j\geq 1}(T_{k,k+j-1})_{*}m_{k}(\tau_{k+j}\geq N_{\#})\leq 1/2$.
  • (ii)
    there exist $\delta_{\#}>0$ and coprime integers $p_{1},\ldots p_{M}$ such that $m_{k}(\tau_{k}=p_{m})\geq\delta_{\#}$ for all $k\geq 1$ and $1\leq m\leq M$.
[…]
Proof.
The proof is similar to in the stationary case [15, Section 4.1]. For completeness, we provide more details in Appendix A. ∎
2.2. Memory loss
We consider a sequence of maps $(T_{n})$ satisfying (NU:1-5), equipped with a sequence of tail bounds $(h^{k})$.
Proposition 2.2.
There exist constants $0<K_{1}<K_{2}$, depending only on $\lambda$ and $K$, such that for each $k\geq 1$ and for each nonnegative measure $\mu$ on $Y_{k}$ with $|\mu|_{{\mathrm{LL}},k}\leq K_{2}$,
$\bigl{|}(T_{k,k+\tau_{k}(a)-1})_{*}(\mu|_{a})\bigr{|}_{{\mathrm{LL}},k+\tau_{k}(a)}\leq K_{1},$
whenever $a\in{\mathcal{P}}_{k}$, $a\subset Y_{k}$. The constants $K_{1}$, $K_{2}$ can be chosen arbitrarily large.
Remark 2.3.
If $\mu$ is a measure supported on $Y_{k}$ and $a\in{\mathcal{P}}_{k}$, then $(T_{k,k+\tau_{k}(a)-1})_{*}(\mu|_{a})$ is supported on $Y_{k+\tau_{k}(a)}$.
Proof of Proposition 2.2.
It is standard, see e.g. [15, Proposition 3.1], that
$\bigl{|}(T_{k,k+\tau_{k}(a)-1})_{*}(\mu|_{a})\bigr{|}_{{\mathrm{LL}},k+\tau_{k}(a)}\leq K+\lambda^{-1}|\mu|_{{\mathrm{LL}},k}.$
We can choose any $K_{2}>(1-\lambda^{-1})^{-1}K$ and $K_{1}=K+\lambda^{-1}K_{2}$. ∎
Fix $K_{1}$, $K_{2}$ as in Proposition 2.2. The notion of regularity defined below is a slight modification of the one considered in [16].
Definition 2.4.
Let $k\geq 1$. A nonnegative measure $\mu$ on $X$ is called regular with respect to $T_{k},T_{k+1},\ldots$ if for every $\ell\geq 1$, setting $a_{k,\ell}=\{x\in X:\tau_{k}(x)=\ell\}$,
$\bigl{|}(T_{k,k+\ell-1})_{*}(\mu|_{a_{k,\ell}})\bigr{|}_{{\mathrm{LL}},k+\ell}\leq K_{1}.$ (2.2)
Given $r:\{0,1,\ldots\}\to[0,\infty)$, we say that $\mu$ has tail bound $r$ with respect to $T_{k},T_{k+1},\ldots$ if for all $n\geq 0$,
$\mu\bigl{(}\{x\in X:\tau_{k}(x)\geq n\}\bigr{)}\leq r(n).$ (2.3)
We say that $\mu$ is regular with tail bound $r$ w.r.t. $T_{k},T_{k+1},\ldots$, if both (2.2) and (2.3) are satisfied.
We gather in the following result some basic properties about regular measures that will be used later.
Proposition 2.5.
Let $k\geq 1$ and $n\geq 0$ be arbitrary.
  • (a)
    The measure $m_{k}$ is regular with tail bound $r=h^{k}$ w.r.t. $T_{k},T_{k+1},\ldots$, and every measure $\mu$ on $Y_{k}$ with $|\mu|_{{\mathrm{LL}},k}\leq K_{2}$ is regular with tail bound $r=\mu(Y_{k})e^{\operatorname{diam}(X)K_{2}}h^{k}$ w.r.t. $T_{k},T_{k+1},\ldots$
  • (b)
    If $\{\mu_{j}\}$ is a finite or countable collection of measures regular w.r.t. $T_{k},T_{k+1},\ldots$, then $\mu=\sum_{j}\mu_{j}$ is regular w.r.t. $T_{k},T_{k+1},\ldots$
  • (c)
    If $\mu$ is a regular measure w.r.t. $T_{k},T_{k+1},\ldots$, then both $(T_{k,k+n-1})_{*}\mu$ and $((T_{k,k+n-1})_{*}\mu)|_{X\setminus Y_{k+n}}$ are regular w.r.t. $T_{k+n},T_{k+n+1},\ldots$ Moreover, if $n\geq 1$,
    $\Bigl{|}\bigl{(}(T_{k,k+n-1})_{*}\mu\bigr{)}\big{|}_{Y_{k+n}}\Bigr{|}_{{\mathrm{LL}},k+n}\leq K_{1}.$
Proof.
(a) The statements about $m_{k}$ follow by (NU:3) and the definition of $h^{k}$, while the statements about $\mu$ follow by Proposition 2.2 combined with the estimate $d\mu/dm_{k}\leq\mu(Y_{k})e^{\operatorname{diam}(X)|\mu|_{{\mathrm{LL}},k}}$.
(b) The statement follows from the fact that if $\nu=\sum_{j}\nu_{j}$, where $\{\nu_{j}\}$ is a countable or finite collection of nonnegative measures on $Y_{k}$, then $|\nu|_{{\mathrm{LL}},k}\leq\sup_{j}|\nu_{j}|_{{\mathrm{LL}},k}$.
(c) It suffices to treat the case $n=1$. Let $\mu$ be a regular measure w.r.t. $T_{k},T_{k+1},\ldots$ We have to verify that $(T_{k})_{*}\mu$ is regular w.r.t. $T_{k+1},T_{k+2},\ldots$ To this end, let $\ell\geq 1$ and let $a_{k,\ell}$ be as in Definition 2.4. Let $\mu^{\prime}=((T_{k})_{*}\mu)|_{Y_{k+1}}$. Then
$(T_{k+1,k+1+\ell-1})_{*}[((T_{k})_{*}\mu)|_{a_{k+1,\ell}}]$ $=(T_{k,k+1+\ell-1})_{*}[\mu|_{a_{k,\ell+1}}]+(T_{k+1,k+1+\ell-1})_{*}[\mu^{\prime}|_{a_{k+1,\ell}}]$
$=\mu_{1}+\mu_{2}.$
Since $\mu$ is regular, $|\mu_{1}|_{{\mathrm{LL}},k+1+\ell}\leq K_{1}$, and since $|\mu^{\prime}|_{{\mathrm{LL}},k+1}\leq K_{1}$, by Proposition 2.2, $|\mu_{2}|_{{\mathrm{LL}},k+1+\ell}\leq K_{1}$ which implies the corresponding bound for the left hand side above, as required. ∎
With this preparation, our main result can be stated as follows:
Theorem 2.6.
Suppose that $\mu$ is a regular probability measure with tail bound $r$ w.r.t. $T_{k},T_{k+1},\ldots$ for some $k\geq 1$. Then there exists a decomposition
$\mu=\sum_{n=1}^{\infty}\alpha_{n}\mu_{n},$
where $\mu_{n}$ are probability measures and $\alpha_{n}$ are nonnegative constants with $\sum_{n\geq 1}\alpha_{n}=1$ such that $(T_{k,k+n-1})_{*}\mu_{n}=m_{k+n}$ for each $n\geq 1$. The sequence $\alpha_{n}$ is fully determined by $K_{1}$, $K_{2}$, the system constants ($\operatorname{diam}X$, $K$, $\lambda$, $n_{0}$, $\delta_{0}$), and the functions $r$ and $h^{j}$, $j\geq k$. In particular, $\alpha_{n}$ does not depend on $\mu$ in any other way. Moreover, suppose there exist constants $0<\beta^{\prime}\leq\beta$ with $\beta>1$, $C_{\beta},C_{\beta}^{\prime}\geq 1$, and a sequence of numbers $\Theta_{j}\in[0,1)$ such that
$h^{j}(n)\leq C_{\beta}(1\vee(n-\Theta_{j}j))^{-\beta}\quad\text{and}\quad r(n)\leq C_{\beta}^{\prime}(1\vee(n-\Theta_{k}k))^{-\beta^{\prime}}$ (2.4)
hold for all $j\geq k$ and $n\geq 1$. Then, if $\sup_{j\geq 1}\Theta_{j}\leq\Theta$ holds for a sufficiently small $\Theta\in(0,1)$ (depending on $\beta,\beta^{\prime}$, and the system constants), we have that for all $n\geq 1$,
$\sum_{j\geq n}\alpha_{j}\leq CC_{\beta}^{\prime}(\Theta_{k}^{*}k+1)^{\beta^{\prime}}n^{-\beta^{\prime}}.$
Here, $\Theta_{k}^{*}=\sup_{j\geq k}\Theta_{j}$ and $C$ is a constant depending only on $n_{0},\Theta,C_{\beta},\beta,\beta^{\prime}$ and $\operatorname{diam}X$, $K$, $\lambda$, $n_{0}$, $\delta_{0}$.
Remark 2.7.
Suppose that $\mu$ and $\nu$ are regular probability measures with the same tail bound $r$ w.r.t. $T_{k},T_{k+1},\ldots$ Under the assumption (2.4), by Theorem 2.6 we obtain the following upper bound on memory loss:
$|(T_{k,k+n-1})_{*}\mu-(T_{k,k+n-1})_{*}\nu|\leq 2\sum_{j>n}\alpha_{j}\leq 2CC_{\beta}^{\prime}(\Theta_{k}^{*}k+1)^{\beta^{\prime}}n^{-\beta^{\prime}}.$
Remark 2.8.
In Section 3, we apply Theorem 2.6 to intermittent maps composed along random ergodic sequences of parameters. In this case, $\Theta_{n}$ is related to the rate of convergence in Birkhoff’s ergodic theorem. For further details, see Corollary 3.3 and Remark 3.6.
3. Applications
3.1. Liverani–Saussol–Vaienti maps
We consider sequences $(T_{n})$ of maps in the family (1.2). First, we prove Theorem 1.1 concerning the rate of memory loss for deterministic compositions, after which we look at applications for random compositions.
Fix a sequence $(T_{k})$ of maps in (1.2) with parameters $\gamma_{k}\in(0,1)$ such that
$\sup_{k\geq 1}\gamma_{k}=\gamma^{*}<1.$
Let $m$ be the Lebesgue measure on $Y=[1/2,1]$ normalized to probability. Denote by $g_{k}:X\to[0,1/2]$ and $h:X\to Y$ respectively the left and right inverse branches of $T_{k}$. Set
$x_{n}(k)=g_{k}\cdots g_{k+n-1}(1)\quad\text{and}\quad y_{n}(k)=hg_{k+1}\cdots g_{k+n-1}(1).$
Then, for all $k\geq 1$ and $n\geq 1$, $T_{k,k+n-1}$ maps both $[x_{n+1}(k),x_{n}(k)]$ and $[y_{n+1}(k),y_{n}(k)]$ bijectively onto $Y$. For each $k\geq 1$, we define a partition of $X$ by
${\mathcal{P}}_{k}=\{[x_{n+1}(k),x_{n}(k)]\>:\>n\geq 1\}\cup\{[y_{n+1}(k),y_{n}(k)]\>:\>n\geq 1\}.$
It is known (see e.g. [1, 2]) that
$x_{n}(k)=O(n^{-1/\gamma_{*}})\quad\text{and}\quad y_{n}(k)-1/2=O(n^{-1/\gamma_{*}}).$ (3.1)
Now, set $m_{k}=m$ and $Y_{k}=Y$ for all $k\geq 1$. It was shown in [16] that (NU:1-5) are satisfied, and that every probability measure $\mu$ with density in ${\mathcal{C}}_{*}(\gamma)$ is regular w.r.t. $T_{k},T_{k+1},\ldots$ for any $k\geq 1$. Thus, Theorem 1.1 follows by combining Theorem 2.6 with the following result.
Proposition 3.1.
Let $\gamma\in(0,1)$. Assume that there exist $a>0$, $\kappa\in(0,1)$, and $N\in{\mathbb{N}}$ such that for all $n\geq N$,
|$\frac{|\{1\leq k\leq n\>:\>\gamma_{k}\leq\gamma\}|}{n}\in[a(1-\kappa),a(1+\kappa)].$ (3.2)
Let $\mu$ be a probability measure with density in ${\mathcal{C}}_{*}(\gamma)$. Then $\mu$ has tail bound $r(n)=O(n^{1-1/\gamma})$ with respect to $T_{1},T_{2},\ldots$ Moreover, for each $k\geq 1$, $h^{k}$ in (2.1) satisfies $h^{k}(\ell)\leq C(1\vee(\ell-\Theta k))^{-1/\gamma}$ where $C$ is a constant independent of $k$, and $\Theta=4\kappa(1-\kappa)^{-1}$.
Proof.
Let $\gamma_{n_{1}},\gamma_{n_{2}},\ldots$ be the sub-sequence of all parameters $\gamma_{n_{j}}$ with $\gamma_{n_{j}}\leq\gamma$, which by (3.2) must be an infinite sequence. Let $\widetilde{T}_{j}=T_{n_{j}}$. It follows by (3.1) that $\mu$ has tail bound $r(\ell)=O(\ell^{1-1/\gamma})$ with respect to the sequence $\widetilde{T}_{1},\widetilde{T}_{2},\ldots$
Next, for $k,\ell\geq 1$, denote
$G_{\ell}(k)$ =#$=\#\{k\leq j\leq k+\ell-1\>:\>\gamma_{j}\leq\gamma\}=\#\bigl{(}\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr{)}.$
Recall that $\tau_{k}$ denotes the first return time to $Y$ w.r.t. $T_{k},T_{k+1},\ldots$ We observe that
$\tau_{k}(x)\geq\ell\implies\widetilde{\tau}_{k}(x)\geq G_{\ell}(k),$
where $\widetilde{\tau}_{k}$ denotes the first return time to $Y$ w.r.t. $T_{n_{m(k)}},T_{n_{m(k)+1}}$, and
$m(k)=\min\{j\in{\mathbb{N}}\>:\>n_{j}\geq k\}.$
Hence, using (3.2) we see that $\mu$ has tail bound $r(\ell)=O(\ell^{1-1/\gamma})$ w.r.t. $T_{1},T_{2},\ldots$:
$\mu(\tau_{1}\geq\ell)\leq\mu(\widetilde{\tau}_{1}\geq G_{\ell}(1))=O(G_{\ell}(1)^{1-1/\gamma})=O((\ell a(1-\kappa))^{1-1/\gamma})=O(\ell^{1-1/\gamma}).$
Moreover,
$m(\tau_{k}\geq\ell)\leq m(\widetilde{\tau}_{k}\geq G_{\ell}(k))=O(G_{\ell}(k)^{-1/\gamma}),$
where (3.1) was used in the last inequality. Finally, whenever $k>N$, it follows from (3.2) that
$G_{\ell}(k)$ $=G_{k+\ell-1}(1)-G_{k-1}(1)\geq(k+\ell-1)a(1-\kappa)-(k-1)a(1+\kappa)$
$\geq\frac{a(1-\kappa)}{2}\biggl{(}\ell-\frac{4\kappa}{1-\kappa}k\biggr{)},$ (3.3)
provided that $\ell\geq 4(1+\kappa)/(1-\kappa)$. ∎
Remark 3.2.
Our method can also be used to treat variations of (1.2) where the right branch contracts near a critical point, as in the following example considered by Cui in [6], which involves two parameters $\gamma\in(0,1)$ and $\beta\geq 1$ (see Figure 1):
$T(x)=\begin{cases}x(1+2^{\gamma}x^{\gamma}),&x\in[0,\tfrac{1}{2}),\\ 2^{\beta}(x-\tfrac{1}{2})^{\beta},&x\in[\tfrac{1}{2},1].\end{cases}$ (3.4)
The map preserves an absolutely continuous probability measure if and only if $\gamma\beta<1$. For sequences $T_{1},T_{2},\ldots$ of maps in (3.4) with parameters $\gamma_{k}\leq\gamma_{*}$ and $\gamma_{k}\leq\beta_{*}$ where $\gamma_{*}\beta_{*}<1$, an application of Theorem 2.6 yields the following estimate provided that (3.2) holds: for any two probability measures $\mu$ and $\mu^{\prime}$ with Hölder continuous densities,
$|(T_{1,n})_{*}(\mu-\mu^{\prime})|=O(n^{-1/\gamma\beta_{*}}).$
3.1.1. Random compositions
Let $\omega=(\omega_{n})$ be a random sequence of parameters sampled from a probability space $(\Omega,{\mathcal{F}},\mathbf{P})=(\Omega_{0}^{{\mathbb{N}}},{\mathcal{E}}^{{\mathbb{N}}},\mathbf{P})$, where $\Omega_{0}\subset(0,1)$ is an interval, ${\mathcal{E}}$ is the Borel sigma-algebra on $\Omega_{0}$, and =${\mathbb{N}}=\{1,2,\ldots\}$. Denote by $T_{n}\circ\cdots\circ T_{1}$ compositions of maps $T_{k}$ along the random sequence $\omega$, where each $T_{k}$ corresponds to a map with parameter $\omega_{k}$. As a consequence of Theorem 1.1, we obtain the following result:
Corollary 3.3.
Let $\gamma\in\Omega_{0}$. Let $\mu,\mu^{\prime}$ and $\nu$ be as in Theorem 1.1.
  • (i)
    Suppose that there exists $b>0$ such that for $\mathbf{P}$-a.e. $\omega\in\Omega$,
    $\lim_{n\to\infty}\frac{1}{n}\sum_{k=0}^{n}\mathbf{1}_{[0,\gamma]}(\omega_{k})=b.$ (3.5)
    Then, for $\mathbf{P}$-a.e. $\omega\in\Omega$, bounds (a) and (b) in Theorem 1.1 hold.
  • (ii)
    Suppose that the following three conditions hold:
    • (A1)
      the shift map $\sigma:\Omega\to\Omega$, $(\sigma\omega)_{n}=\omega_{n+1}$, preserves $\mathbf{P}$,
    • (A2)
      $(\sigma,\mathbf{P})$ is ergodic, and
    • (A3)
      $\mathbf{P}(\omega_{1}\leq\gamma)>0$.
    Then, for $\mathbf{P}$-a.e. $\omega\in\Omega$, bounds (a) and (b) in Theorem 1.1 hold.
Proof.
Clearly (3.5) implies (1.4), so (i) follows by Theorem 1.1. For (ii), let $\varphi=\mathbf{1}_{[0,\gamma]}\circ\pi_{1}$ where $\pi_{1}(\omega)=\omega_{1}$ projects onto the first coordinate. By Birkhoff’s ergodic theorem, we have
$\lim_{n\to\infty}\frac{1}{n}\sum_{k=0}^{n}\mathbf{1}_{[0,\gamma]}(\omega_{k})=\lim_{n\to\infty}n^{-1}\sum_{k=0}^{n}\varphi\circ\sigma^{k}=\mathbf{P}(\omega_{1}\leq\gamma)>0$
for $\mathbf{P}$-a.e. $\omega\in\Omega$. Hence, (3.5) holds with $b=\mathbf{P}(\omega_{1}\leq\gamma)$. ∎
Remark 3.4.
The constants in the memory loss bounds of Corollary 3.3 depend on $\omega$ through $N=N(\omega)$ in Theorem 1.1.
Remark 3.5.
In the setting of Corollary 3.3-(ii), if $\mathbf{P}(\omega_{1}\leq\gamma)>0$ holds for all $\gamma>0$, then we obtain $|(T_{1,n})_{*}(\mu-\mu^{\prime})|=O(n^{-1/\gamma})$ for all $\gamma>0$, i.e. memory loss occurs at a superpolynomial rate.
Remark 3.6.
In the setting of Corollary 3.3(i), suppose that for $\mathbf{P}$-a.e. $\omega\in\Omega$,
$\biggl{|}\frac{1}{n}\sum_{k=0}^{n-1}\mathbf{1}_{[0,\gamma]}(\omega_{k})-b\biggr{|}\leq\Theta_{n},$
where $\Theta_{n}=\Theta_{n}(\omega)\in{\mathbb{R}}_{+}$ and $\lim_{n\to\infty}\Theta_{n}=0$. Then, similar to (3.1), we find that for sufficiently large $n$ and all $j\geq 1$,
$h^{j}(n)\leq C\biggl{(}1\vee(n-C^{\prime}\Theta_{j}^{*}j)\biggr{)}^{-1/\gamma}$
holds for some constants $C$ and $C^{\prime}$ (depending on $\omega$), where $\Theta_{j}^{*}=\sup_{\ell\geq j}\Theta_{\ell}$. By Theorem 2.6, it follows that for any two probability measures $\mu$ and $\mu^{\prime}$ with Hölder continuous densities,
$|(T_{k,k+n-1})_{*}\mu-(T_{k,k+n-1})_{*}\mu^{\prime}|=O((\Theta_{k}^{*}k+1)^{1/\gamma}n^{-1/\gamma})$
holds for $\mathbf{P}$-a.e. $\omega\in\Omega$, where the constant is independent of $k$. For example, if $\Theta_{j}=O(j^{-\psi})$ for some $\psi\in(0,1]$, then
$|(T_{n,n+\lceil n^{a}\rceil-1})_{*}\mu-(T_{n,n+\lceil n^{a}\rceil-1})_{*}\mu^{\prime}|=O(n^{(1-\psi-a)/\gamma})$
for $\mathbf{P}$-a.e. $\omega\in\Omega$.
In situations where the shift map $\sigma$ does not preserve the measure $\mathbf{P}$, alternative conditions can be formulated to ensure that (3.5) is satisfied. The following result provides an example in the case of asymptotically mean stationary strongly mixing noise.
Proposition 3.7.
Let $\gamma\in\Omega_{0}$. Assume the following:
  • (B1)
    There exists a probability measure $\bar{\mathbf{P}}$ such that for any bounded measurable function $g:\Omega\to{\mathbb{R}}$,
[…]
  • (B2)
    The strong mixing coefficients
    $\alpha(n)=\sup_{i\geq 1}\alpha({\mathcal{F}}_{1}^{i},{\mathcal{F}}_{i+n}^{\infty})$
    satisfy
    $\alpha(n)=O(\log^{-\beta}(n))\quad\text{for some $\beta>1$.}$
    Here, ${\mathcal{F}}_{1}^{i}$ is the sigma-algebra on $\Omega$ generated by the projections $\pi_{1},\ldots,\pi_{i}$; ${\mathcal{F}}_{i+n}^{\infty}$ is the sigma-algebra generated by $\pi_{i+n}\pi_{i+n+1},\ldots$, with $\pi_{k}(\omega)=\omega_{k}$; and
    $\alpha({\mathcal{F}}_{1}^{i},{\mathcal{F}}_{j}^{\infty})=\sup_{A\in{\mathcal{F}}_{1}^{i},\,B\in{\mathcal{F}}_{j}^{\infty}}|\mathbf{P}(A\cap B)-\mathbf{P}(A)\mathbf{P}(B)|.$
  • (B3)
    $\bar{\mathbf{P}}(\omega_{1}\leq\gamma)>0$.
Then, (3.5) holds.
Proof.
Define $\varphi$ as in the proof of Corollary 3.3. By (B3), it is enough to show that, for $\mathbf{P}$-a.e. $\omega\in\Omega$,
$\lim_{n\to\infty}n^{-1}\sum_{k=0}^{n-1}\varphi\circ\sigma^{k}=\bar{\mathbf{P}}(\omega_{1}\leq\gamma).$ (3.6)
To this end, we decompose $n^{-1}\sum_{k=0}^{n-1}\varphi\circ\sigma^{k}-\bar{\mathbf{P}}(\omega_{1}\leq\gamma)=I+II$, where
[…]
and
[…]
By (B1), $I\to 0$ as $n\to\infty$. By (B2), $(\xi_{n})$ is a uniformly bounded sequence of $\alpha$-mixing random variables whose $\alpha$-mixing coefficients decay at the rate $O(\log^{-\beta}(n))$ with $\beta>1$. It follows from [26, Corollary 1] that $II\to 0$ as $n\to\infty$, for $\mathbf{P}$-a.e. $\omega\in\Omega$. Thus, we obtain (3.6). ∎
3.2. Pikovsky maps
For $\gamma>1$, let $T:X\to X$ be a map on $X=[-1,1]$ defined implicitly by the relations
$x=\begin{cases}\frac{1}{2\gamma}(1+T(x))^{\gamma},&0\leq x\leq\frac{1}{2\gamma},\\ T(x)+\frac{1}{2\gamma}(1-T(x))^{\gamma},&\frac{1}{2\gamma}\leq x\leq 1,\end{cases}$ (3.7)
and by letting $T(x)=-T(-x)$ for $x\in[-1,0]$. The graph of $T$ is illustrated in Figure 1. The map has neutral fixed points at $x=1,-1$, while at $x=0$ its derivative becomes infinite. There is an interplay between the neutrality of the fixed points and the degree of the singularity, which is controlled by the parameter $\gamma$:
$T^{\prime}(x)\approx C_{\gamma}x^{1/\gamma-1}-1\quad\text{as $x\searrow 0$},$
$T^{\prime}(x)\approx 1+\frac{1}{2}(1-x)^{\gamma-1}\quad\text{as $x\nearrow 1$}.$
Note that $T$ preserves the Lebesgue measure, and that $T$ reduces to the angle doubling map at the limit $\gamma\searrow 1$. By [4, Proposition 5], $T$ is polynomially mixing with rate $n^{-1/(\gamma-1)}$ for Hölder continuous observables. This result was recently extended in [20] to i.i.d. random compositions of maps in the family (3.7).
Let $(T_{k})$ be a sequence of maps in (3.7) with parameters $\gamma_{1},\gamma_{2},\ldots$ We assume that, for all $k\geq 1$,
$\gamma_{k}\in[\gamma_{-},\gamma_{+}]\quad\text{where $1<\gamma_{-}<\gamma_{+}<3$.}$
By applying Theorem 2.6 in combination with certain technical estimates from [4, 20], we obtain the following bound on memory loss for deterministic compositions:
Theorem 3.8.
Let $\gamma_{-}\leq\gamma\leq\gamma_{+}$. There exists $\kappa_{0}\in(0,1)$ such that the following holds. Assume that for some $a>0$, $\kappa\leq\kappa_{0}$, and $N\in{\mathbb{N}}$,
|$\frac{|\{1\leq k\leq n\>:\>\gamma_{k}\leq\gamma\}|}{n}\in[a(1-\kappa),a(1+\kappa)]$ (3.8)
holds for all $n\geq N$. Let $\mu$ and $\mu^{\prime}$ be probability measures on $X$ with Hölder continuous densities. Then
$|(T_{1,n})_{*}(\mu-\mu^{\prime})|=O(n^{-\frac{1}{\gamma-1}}).$ (3.9)
Remark 3.9.
A concrete expression for $\kappa_{0}$ depending on dynamical constants can be extracted from the proofs of Propositions 3.12 and 4.6.
Before proving Theorem 3.8, we state a corollary for random compositions. Let $(\omega_{n})$ be a random sequence of parameters sampled from a probability space $(\Omega,{\mathcal{F}},\mathbf{P})=(\Omega_{0}^{{\mathbb{N}}},{\mathcal{E}}^{{\mathbb{N}}},\mathbf{P})$, where $\Omega_{0}=[\gamma_{-},\gamma_{+}]$ with $31<\gamma_{-}<\gamma_{+}<$ 3, and ${\mathcal{E}}$ is the Borel $\sigma$-algebra on $\Omega_{0}$. As in Section 3.1.1, denote by $T_{n}\circ\cdots\circ T_{1}$ compositions of maps $T_{k}$ from (3.7) along the random sequence $\omega$, where each $T_{k}$ corresponds to a map with parameter $\omega_{k}$.
Corollary 3.10.
Let $\gamma\in\Omega_{0}$. Assume the following:
  • (A1)
    the law $\mathbf{P}$ is preserved by the shift map $\sigma$,
  • (A2)
    $(\sigma,\mathbf{P})$ is ergodic, and
  • (A3)
    $\mathbf{P}(\omega_{1}\leq\gamma)>0$.
Let $\mu$ and $\nu$ be probability measures on $X$ with Hölder continuous densities. Then
$|(T_{1,n})_{*}(\mu-\nu)|=O(n^{-\frac{1}{\gamma-1}}).$ (3.10)
Proof.
The result follows by combining Theorem 3.8 with Birkhoff’s ergodic theorem, as in the proof of Corollary 3.3. ∎
3.2.1. Proof of Theorem 3.8
Proving Theorem 3.8 amounts to verifying that nonstationary compositions of Pikovsky maps (3.7) fit the abstract framework in Section 2, and demonstrating that under (3.8), the tails of return times decay at the correct order in view of (3.9). Both of these properties are essentially a consequence of certain estimates established in [4, 20]. In particular, (NU:3) follows from the distortion bounds in [20, Lemma 2.11]. Further details are provided below.
Let $(T_{k})$ be a sequence of maps in the family (3.7) with parameters $\gamma_{k}\in[\gamma_{-},\gamma_{+}]$, where $31<\gamma_{-}<\gamma_{+}<$ 3. Let $X_{-}=(-0,1)$ and $X_{+}=(0,1)$. Let $\lambda$ be the Lebesgue measure on $X$ normalized to probability. Denote by $g_{k,-}:(-1,1)\to X_{-}$ and $g_{k,-}:(-1,1)\to X_{+}$ the left and right inverse branches of $T_{k}$, respectively.
To define ${\mathcal{P}}_{k}$ in Section 2, we recall from [4, 20] the definitions of certain partitions associated with the sequence $(T_{k})$. First, for each $n\geq 1$ and $k\geq 1$, let
$\Delta_{n}^{-}(k)=g_{k,-}(\Delta_{n-1}^{-}(k+1))\quad\text{and}\quad\Delta_{n}^{+}(k)=g_{k,+}(\Delta_{n-1}^{+}(k+1)),$
where
$\Delta_{0}^{-}(k)=(g_{k,-}(0),0)=g_{k,-}(X_{+})\quad\forall k\geq 1,$
and
$\Delta_{0}^{+}(k)=(0,g_{k,+}(0))=g_{k,+}(X_{-})\quad\forall k\geq 1.$
Then $\{\Delta_{n}^{-}(k)\}_{n\geq 1}$ is a (mod $\lambda$) partition of $X_{-}$, $\{\Delta_{n}^{+}(k)\}_{n\geq 1}$ is a (mod $\lambda$) partition of $X_{+}$, and $T_{k}$ maps $\Delta_{n}^{\pm}(k)$ bijectively onto $\Delta_{n-1}^{\pm}(k+1)$. Moreover, $\Delta^{+}_{\ell}(k)=(x_{\ell}^{+}(k),x_{\ell+1}^{+}(k))$ and $\Delta^{-}_{\ell}(k)=(x_{\ell+1}^{-}(k),x_{\ell}^{-}(k))$, where $x_{\ell}^{+}(k)=g_{k,+}\cdots g_{k+\ell-1,+}(0)$.
Next, for each $k,n\geq 1$ define
$\delta_{n}^{-}(k)=T_{k}^{-1}(\Delta_{n-1}^{+}(k+1))\cap\Delta_{0}^{-}(k)\quad\text{and}\quad\delta_{n}^{+}(k)=T_{k}^{-1}(\Delta_{n-1}^{-}(k+1))\cap\Delta_{0}^{+}(k).$
Then $\{\delta_{n}^{-}(k)\}_{n\geq 1}$ and $\{\delta_{n}^{+}(k)\}_{n\geq 1}$ are (mod $\lambda$) partitions of $\Delta_{0}^{-}(k)$ and $\Delta_{0}^{+}(k)$, respectively. Note that the following maps are all bijective:
$T_{k}:\delta_{n}^{-}(k)\to\Delta_{n-1}^{+}(k+1),\quad T_{k}:\delta_{n}^{+}(k)\to\Delta_{n-1}^{-}(k+1),$
$T_{k,k+n-1}:\delta_{n}^{-}(k)\to\Delta_{0}^{+}(k+n),\quad T_{k,k+n-1}:\delta_{n}^{+}(k)\to\Delta_{0}^{-}(k+n).$
Set
$Y_{k}=\Delta_{0}^{-}(k)\cup\Delta_{0}^{+}(k).$ (3.11)
Let $m_{k}$ be the Lebesgue measure on $Y_{k}$ normalized to probability. For each $k\geq 1$, we define a partition ${\mathcal{P}}_{k}$ of $X$ as follows:
${\mathcal{P}}_{k}=\{\delta_{n}^{-}(k)\cup\delta_{n}^{+}(k)\}_{n\geq 1}\cup\{\Delta_{n}^{-}(k)\cup\Delta_{n}^{+}(k)\}_{n\geq 1}.$
By definition, $T_{k,k+n-1}$ maps $\delta_{n}(k):=\delta_{n}^{-}(k)\cup\delta_{n}^{+}(k)$ bijectively onto $Y_{k+n}$, and $T_{k,k+n-1}$ maps $\Delta_{n}(k):=\Delta_{n}^{-}(k)\cup\Delta_{n}^{+}(k)$ bijectively onto $Y_{k+n}$. For $x\in X$, let
$\tau_{k}(x)=\inf\{n\geq 1\>:\>T_{k,k+n-1}(x)\in Y_{k+n}\}.$
Then $\tau_{k}$ is constant on the partition elements of ${\mathcal{P}}_{k}$, so that (NU:2) is satisfied. By inspecting the definition of $T$ in (3.7), we see that $|F_{a}(y)-F_{a}(y^{\prime})|\geq\Lambda|y-y^{\prime}|>0$ holds for all $y,y^{\prime}\in a\subset Y_{k}$ whenever $a\in{\mathcal{P}}_{k}$ where $F_{a}=T_{k,k+\tau_{k}(a)-1}$ and $\Lambda=\Lambda(\gamma_{-},\gamma_{+})>1$. Then (NU:3) follows by [20, Lemma 2.11] provided that $31<\gamma_{-}<\gamma_{+}<$ 3:
Proposition 3.11.
Assume that $31<\gamma_{-}<\gamma_{+}<$ 3. Then there exists $K>0$ such that for any $k\geq 1$ and $a\subset Y_{k}$, $a\in{\mathcal{P}}_{k}$,
$\zeta=\frac{d(F_{a})_{*}(m_{k}|_{a})}{dm_{k+\tau_{k}(a)}}\quad\text{satisfies}\quad|\zeta|_{{\mathrm{LL}},k+\tau_{k}(a)}\leq K.$
Proof.
Suppose that $a=\delta_{n}(k)$. Let $x,x^{\prime}\in Y_{k+n}$. Then
$\zeta(x)=\frac{\lambda(Y_{k+\tau_{k}(a)})}{\lambda(Y_{k})}\frac{1}{T_{k,k+n-1}^{\prime}(z)},$
where $z=(T_{k,k+n-1}|_{\delta_{n}(k)})^{-1}(x)$. Let $z^{\prime}=(T_{k,k+n-1}|_{\delta_{n}(k)})^{-1}(x^{\prime})$. If $z,z^{\prime}\in\delta_{n}^{+}(k)$ or $z,z^{\prime}\in\delta_{n}^{-}(k)$, then by [20, Lemma 2.11],
$|\log\zeta(x)-\log\zeta(x^{\prime})|\leq K|T_{k,k+n-1}(z)-T_{k,k+n-1}(z^{\prime})|=K|x-x^{\prime}|.$
Next, suppose that $z$ and $z^{\prime}$ belong to different smoothness components, say $z\in\delta_{n}^{-}(k)$ and $z^{\prime}\in\delta_{n}^{+}(k)$. In this case we can exploit the trick described in [4, Remark 1]. Namely, by symmetry of $T_{k}$ we have $-z^{\prime}\in\delta_{n}^{-}(k)$ and $T_{j}^{\prime}(y)=T_{j}^{\prime}(-x)$. Since each $T_{j}$ is an odd function,
$|\log\zeta(x)-\log\zeta(x^{\prime})|$ $=|\log T_{k,k+n-1}^{\prime}(z)-\log T_{k,k+n-1}^{\prime}(-z^{\prime})|$
$\leq K|T_{k,k+n-1}(z)-T_{k,k+n-1}(-z^{\prime})|=K|x+x^{\prime}|\leq K(x-x^{\prime}).$
Assumption (NU:4) is trivially true, since all maps in (3.7) are expanding. For (NU:5), since each of the maps $T_{n}$ preserves $\lambda$, with $N_{\#}\geq 1$, we have
$(T_{k,k+j-1})_{*}m_{k}(\tau_{k+j}\geq N_{\#})\leq\frac{(T_{k,k+j-1})_{*}\lambda(\tau_{k+j}\geq N_{\#})}{\lambda(Y_{k})}=\frac{\lambda(\tau_{k+j}\geq N_{\#})}{\lambda(Y_{k})}\leq CN_{\#}^{-1/(\gamma_{+}-1)},$
where the last inequality follows from (3.12) below, and $C$ is a constant determined by $\gamma_{-},\gamma_{+}$. Moreover, we have $m_{k}(\tau_{k}=1)\geq m_{k}(\delta_{1}^{-}(k))\geq\delta_{\#}$ for some constant $\delta_{\#}>0$ determined by $\gamma_{-},\gamma_{+}$. Thus, (NU:5) holds by Proposition 2.1.
Proposition 3.12.
Assume (3.8). Then for each $k\geq 1$, $h^{k}$ in (2.1) satisfies $h^{k}(\ell)=O((1\vee(\ell-\Theta k))^{-1/(\gamma-1)})$ where the constant is independent of $k$, and $\Theta=c_{0}\kappa(1-\kappa)^{-1}$ for some absolute constant $c_{0}>0$. Moreover, for any probability measure $\mu$ with bounded density, $\mu(\tau_{1}\geq n)=O(n^{-1/(\gamma-1)})$.
Proof.
First, consider a sequence $(T_{j})$ of maps in (3.8) with parameters $\gamma_{j}\in[\gamma_{-},\gamma]$ where $\gamma$ is as in (3.7). Then
$\lambda(\tau_{k}\geq n)=\lambda\biggl{(}\bigcup_{\ell\geq n}\Delta_{\ell}(k)\biggr{)}+\lambda\biggl{(}\bigcup_{\ell\geq n}\delta_{\ell}(k)\biggr{)}.$
Denoting by $x_{\ell}^{+}(\gamma)$ the preimages corresponding to the stationary sequence $\gamma_{1}=\gamma_{2}=\cdots=\gamma$, we have
$1-x_{\ell}^{+}(k)\leq 1-x_{\ell}^{+}(\gamma)\leq Cn^{-\frac{1}{\gamma-1}},$
where the last inequality is due to [4]. Hence, $\lambda(\cup_{\ell\geq n}\Delta_{\ell}^{+}(k))=\lambda((x_{n-1}(k)^{+},1))=O(n^{-1/(\gamma-1)})$. The same estimate holds with $\Delta_{\ell}^{-}(k)$ in place of $\Delta_{\ell}^{+}(k)$, leading to $\lambda(\cup_{\ell\geq n}\Delta_{\ell}(k))=O(n^{-1/(\gamma-1)})$. On the other hand, by [20, Lemma 2.4],
$\lambda\biggl{(}\bigcup_{\ell\geq n}\delta_{\ell}^{\pm}(k)\biggr{)}=O\biggl{(}n^{-\frac{\gamma_{-}}{\gamma-1}}\biggr{)},$
where the constant is determined by $\gamma_{-},\gamma$. Hence,
$\lambda(\tau_{k}\geq n)=O(n^{-\frac{1}{\gamma-1}}).$ (3.12)
Next, let $(T_{j})$ be a sequence of maps as in Theorem 3.8. Similar to the proof of Theorem 1.1, we let $\widetilde{T}_{j}=T_{n_{j}}$ where $\gamma_{n_{1}},\gamma_{n_{2}},\ldots$ is the subsequence of all parameters $\gamma_{n_{j}}$ with $\gamma_{n_{j}}\leq\gamma$ and, for each $k,\ell\geq 1$, define
$G_{\ell}(k)$ =#$=\#\{k\leq j\leq k+\ell-1\>:\>\gamma_{j}\leq\gamma\}=\#\bigl{(}\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr{)}.$
Further, let
$\widetilde{\tau}_{k}(\ell)=\inf\{\ell\geq 1\>:\>\widetilde{T}_{m(k),m(k)+\ell-1}(x)\in Y_{n_{m(k)+\ell}}\},\quad m(k)=\min\{j\geq 1\>:\>n_{j}\geq k\},$
where $\widetilde{T}_{m(k),m(k)+\ell-1}=\widetilde{T}_{m(k)}\circ\cdots\circ\widetilde{T}_{m(k)+\ell-1}$. By the definitions of $\Delta_{n}(k)$ and $\delta_{n}(k)$, we have
$\bigcup_{j\geq\ell}\Delta_{j}(k)\subset\{\widetilde{\tau}_{k}\geq G_{\ell}(k)\}\quad\text{and}\quad\delta_{j}(k)\subset T_{k}^{-1}\Delta_{j-1}(k+1).$
Hence, using (3.12) we obtain
$\mu(\tau_{k}\geq\ell)$ $\leq\mu(\widetilde{\tau}_{k}\geq G_{\ell}(k))+(T_{k})_{*}\mu(\cup_{j\geq\ell}\Delta_{j-1}(k+1))$
$\leq\mu(\widetilde{\tau}_{k}\geq G_{\ell}(k))+(T_{k})_{*}\mu(\widetilde{\tau}_{k+1}\geq G_{\ell-1}(k+1))$
[…]
$\leq C^{\prime}G_{\ell}(k)^{-\frac{1}{\gamma-1}}+C^{\prime}G_{\ell-1}(k+1)^{-\frac{1}{\gamma-1}},$
where $C^{\prime}>0$ is a constant depending on $\gamma,\gamma_{-},\gamma_{+}$ and $\mu$. Finally, as in (3.1), by invoking (3.8) we see that
$G_{\ell}(k)\geq\frac{a(1-\kappa)}{2}\biggl{(}\ell-\frac{c_{0}\kappa}{1-\kappa}k\biggr{)},$
whenever $k\geq N$ and $\ell$ is large enough (depending on $\kappa$), where $c_{0}>0$ is an absolute constant. ∎
To complete the proof of Theorem 3.8, it remains to verify that measures with Hölder continuous densities are regular.
Proposition 3.13.
Let $\mu$ be a probability measure on $X$ with a positive Hölder continuous density. Then, for sufficiently large $K_{1}$ in the definition of regularity, $\mu$ is regular w.r.t. $T_{k},T_{k+1},\ldots$ for all $k\geq 1$.
Proof.
Let $a\in{\mathcal{P}}_{k}$. Denote by $\rho$ the density of $\mu$. First, assume that $\rho$ is Lipschitz continuous with $\inf\rho=b>0$. Then, with $F_{a}$ as in (NU:3),
$\rho_{a}(y):=\frac{d(F_{a})_{*}(\mu|_{a})}{dm_{k+\tau_{k}(a)}}(y)\propto\frac{\rho(y_{a})}{F_{a}^{\prime}(y_{a})},$
where $y_{a}=F_{a}^{-1}(y)$. Hence, for $y,z\in Y_{k+\tau_{k}(a)}$, we have
$|\log\rho_{a}(y)-\log\rho_{a}(y^{\prime})|\leq I+II,$
where
$I=|\log\rho(y_{a})-\log\rho(z_{a})|\qquad\text{and}\qquad II=|\log F_{a}^{\prime}(y_{a})-\log F_{a}^{\prime}(z_{a})|.$
Since $F_{a}$ is distance expanding on $a$,
$I\leq b^{-1}|\rho(y_{a})-\rho(z_{a})|\leq b^{-1}L|y_{a}-z_{a}|\leq b^{-1}L|y-z|,$
where $L$ is the Lipschitz constant of $\rho$.
By Proposition 3.11, $II\leq K|y-z|$ holds if $a\subset Y_{k}$. Then suppose that $a\subset X\setminus Y_{k}$, say $a=\Delta_{\ell}(k)$. If $z_{a},y_{a}\in\Delta_{\ell}^{+}(k)$, we have $z_{a}=T_{k-1}(z_{a}^{\prime})$ and $y_{a}=T_{k-1}(y_{a}^{\prime})$ for some $z_{a}^{\prime},y_{a}^{\prime}\in\delta_{\ell+1}^{-}(k-1)$. (In the case $k=1$, we let $T_{0}$ be the map with parameter $\gamma_{-}$). It follows by [20, Lemma 2.11] that
$II=|\log T_{k,k+\ell-1}^{\prime}\circ T_{k-1}(y_{a}^{\prime})-\log T_{k,k+\ell-1}^{\prime}\circ T_{k-1}(z_{a}^{\prime})|\leq K|y-z|.$
The same estimate holds if $z_{a},y_{a}\in\Delta_{\ell}^{-}(k)$. If $z_{a}$ and $y_{a}$ belong to different components of $\Delta_{\ell}(k)$, we can repeat the argument from the proof of Proposition 3.11 to again obtain $II\leq K|y-z|$. We conclude that $\rho$ is regular w.r.t. $T_{k},T_{k+1},\ldots$
Finally, if $\rho$ is Hölder continuous with exponent $\eta\in(0,1]$, then $\mu$ is regular with respect to the metric $d_{\eta}(x,y)=|x-y|^{\eta}$. Therefore, we can deal with this case by simply changing the metric from $d(x,y)=|x-y|$ to $d_{\eta}$ in the definitions of Section 2.1. ∎
By Propositions 3.12 and 3.13, the conclusion of Theorem 3.8 holds if the densities of $\mu$ and $\mu^{\prime}$ are positive. Finally, the observation that
$(T_{1,n})_{*}(\mu-\mu^{\prime})=2(T_{1,n})_{*}(\mu_{1}-\mu^{\prime}_{1})$ (3.13)
where
$\mu_{1}=\frac{\mu+1}{2}\quad\text{and}\quad\mu^{\prime}_{1}=\frac{\mu^{\prime}+1}{2}$
are measures with positive Hölder densities, can be used to remove the restriction that the densities of $\mu$ and $\mu^{\prime}$ are positive.
3.3. Grossmann–Horner maps
We follow the representation in [20]. Let $T:X\to X$ be a map on $X=[-1,1]$ with graph as in Figure 1, having two surjective and symmetric branches on the subintervals $[-1,0]$ and $[0,1]$, such that $T|_{X\setminus\{0\}}\in C^{1}$ and $T|_{X\setminus\{0,-1,1\}}\in C^{2}$, defined locally by
$T(x)=\begin{cases}1-b|x|^{\eta}+o(|x|^{\eta}),&\text{if $x\in U_{0}(\gamma,\eta)$},\\ -x+a|x-1|^{\gamma}+o(|x-1|^{\gamma}),&\text{if $x\in U_{1_{-}}(\gamma,\eta)$},\\ x+a|x+1|^{\gamma}+o(|x+1|^{\gamma}),&\text{if $x\in U_{-1_{+}}(\gamma,\eta)$}.\end{cases}$ (3.14)
Here, $a,b>0$, $\eta>0$ and $\gamma>1$ are parameters, $U_{0}(\gamma,\eta)$ is a small neighborhood of 00, and $U_{1_{-}}(\gamma,\eta),U_{-1_{+}}(\gamma,\eta)$ are small neighborhoods of $1,-1$, respectively. For $i=1,2$, the derivatives of $T$ are assumed to satisfy
$|T^{(i)}(x)|$ $=C_{i,0}|x|^{\eta-i}+o(|x|^{\eta-i}),\quad\text{if $x\in U_{0}(\gamma,\eta)$},$
$T^{(i)}(x)$ $=2-i+C_{i,1}|x-1|^{\gamma-i}+o(|x-1|^{\gamma-i})\quad\text{if $x\in U_{1_{-}}(\gamma,\eta)$},$ (3.15)
$T^{(i)}(x)$ $=i-2+C_{i,-1}|x+1|^{\gamma-i}+o(|x+1|^{\gamma-i})\quad\text{if $x\in U_{-1_{+}}(\gamma,\eta)$},$
where $C_{i,j}>0$ are constants. Further, assume:
  • $|T^{\prime}(x)|>1$ for all $x\neq-1,1$,
  • $T$ is strictly monotonous and convex on the two intervals $[-1,0]$ and $[0,1]$, and
  • the error terms in (3.14) and (3.3) are uniformly bounded.
It was shown in [4] that, for $0<\eta<\min\{1/(\gamma-1),1\}$, there exists an absolutely continuous invariant probability measure $\mu$ that mixes polynomially fast for Hölder continuous observables with rate $n^{1-1/\eta(\gamma-1)}$. A nonautonomous variant of this result was obtained in [20], who considered random i.i.d. compositions of maps in (3.14). The memory loss bound below follows by applying Theorem 2.6 in combination with results in [20].
Fix $0<a_{-}<a_{+}<\infty$, $0<b_{-}<b_{+}<\infty$, $1<\gamma_{-}<\gamma_{+}\leq 2$, $10<\eta_{-}<\eta_{+}<$ 1. Let $T_{1},T_{2},\ldots$ be a sequence of maps in (3.14), such that the parameters $a_{i},b_{i},\eta_{i},\gamma_{i}$ associated to $T_{i}$ satisfy
$\eta_{-}\leq\eta_{i}\leq\eta_{+},\quad\gamma_{-}\leq\gamma_{i}\leq\gamma_{+},\quad a_{-}\leq a_{i}\leq a_{+},\quad b_{-}\leq b_{i}\leq b_{+}.$ (3.16)
We also assume that the constants $C_{i,j}$ appearing in (3.3) satisfy the uniform bounds $0<C_{-}\leq C_{i,j}\leq C_{+}<\infty$.
Theorem 3.14.
Let $\gamma\leq\gamma_{+}$. There exists $\kappa_{0}\in(0,1)$ such that the following holds. Assume that for some $a>0$, $\kappa\leq\kappa_{0}$, and $N\in{\mathbb{N}}$,
|$\frac{|\{1\leq k\leq n\>:\>\gamma_{k}\leq\gamma\}|}{n}\in[a(1-\kappa),a(1+\kappa)]$ (3.17)
holds for all $n\geq N$. Let $\mu$ and $\mu^{\prime}$ be probability measures on $X$ with Hölder continuous densities. Then,
$|(T_{1,n})_{*}(\mu-\mu^{\prime})|=O(n^{-1/(\gamma-1)}).$
Remark 3.15.
A concrete expression for $\kappa_{0}$ depending on dynamical constants can be extracted from the proofs of Propositions 3.19 and 4.6.
A bound analogous to (3.10) can be derived by applying Theorem 3.14 in the case of random ergodic compositions of maps in (3.14). We omit the exact statement to avoid repetition.
3.3.1. Proof of Theorem 3.14
Following the representation in [20], we start by describing several partitions associated with nonstationary compositions $T_{n}\circ\cdots\circ T_{1}$ of maps in (3.14). Let $X=[-1,1]$ and, as in Section 3.2.1, denote by $g_{k,-}:(-1,1)\to X_{-}$ and $g_{k,+}:(-1,1)\to X_{+}$ the left and right inverse branches of $T_{k}$, respectively. Here $X_{-}=(-1,0)$ and $X_{+}=(0,1)$. Denote by $\lambda$ the Lebesgue measure on $X$ normalized to probability.
For each $n\geq 1$ and $k\geq 1$, let
$\Delta_{n}^{-}(k)=g_{k,-}(\Delta_{n-1}^{-}(k+1))\quad\text{and}\quad\Delta_{n}^{+}(k)=g_{k,+}(\Delta_{n-1}^{-}(k+1)),$
where
$\Delta_{0}^{-}(k)=(g_{k,-}(0),0)=g_{k,-}(X_{+})\quad\text{and}\quad\Delta_{0}^{+}(k)=(0,g_{k,+}(0))=g_{k,+}(X_{+}).$
Then $\{\Delta_{n}^{-}(k)\}_{n\geq 1}$ and $\{\Delta_{n}^{+}(k)\}_{n\geq 1}$ are (mod $\lambda$) partitions of $X_{-}$ and $X_{+}$, respectively. If $n=0$, $T_{k}$ maps $\Delta_{n}^{\pm}(k)$ bijectively onto $X_{+}$ and, if $n\geq 1$, $T_{k}$ maps $\Delta_{n}^{\pm}(k)$ bijectively onto $\Delta_{n-1}^{-}(k+1)$.
For $n\geq 1$, let
$\delta_{n}^{-}(k)=T_{k}^{-1}(\Delta_{n-1}^{+}(k+1))\cap\Delta_{0}^{-}(k)\quad\text{and}\quad\delta_{n}^{+}(k)=T_{k}^{-1}(\Delta_{n-1}^{+}(k+1))\cap\Delta_{0}^{+}(k).$
Then $\{\delta_{n}^{-}(k)\}_{n\geq 1}$ and $\{\delta_{n}^{+}(k)\}_{n\geq 1}$ are (mod $\lambda$) partitions of $\Delta_{0}^{-}(k)$ and $\Delta_{0}^{+}(k)$, respectively. If $n=1$, $T_{k}$ maps $\delta_{n}^{\pm}(k)$ onto $\Delta_{0}^{+}(k+1)$ and if $n\geq 2$, $T_{k,k+n-1}$ maps $\delta_{n}^{\pm}(k)$ bijectively onto $\Delta_{0}^{-}(k+n)$.
Finally, define (mod $\lambda$) partitions of $\delta_{1}^{-}(k)$ and $\delta_{1}^{+}(k)$ respectively by
$\delta_{1,n}^{-}(k)=T_{k}^{-1}(\delta_{n}^{+}(k+1))\cap\Delta_{0}^{-}(k),\quad n\geq 1,$
and
$\delta_{1,n}^{+}(k)=T_{k}^{-1}(\delta_{n}^{+}(k+1))\cap\Delta_{0}^{+}(k),\quad n\geq 1.$
Then $T_{k,k+n}$ maps both $\delta_{1,n}^{-}(k)$ and $\delta_{1,n}^{+}(k)$ bijectively onto $\Delta_{0}^{-}(k+n+1)$.
Let $Y_{k}=\Delta_{0}^{-}(k)$, and let $m_{k}$ be the Lebesgue measure on $Y_{k}$ normalized to probability. For each $k\geq 1$, define a partition of $X$ by
${\mathcal{P}}_{k}=\{\Delta_{n}^{-}(k),\Delta_{n}^{+}(k)\}_{n\geq 1}\cup\{\delta_{n}^{-}(k),\delta_{n}^{+}(k)\}_{n\geq 2}\cup\{\delta_{1,n}^{-}(k),\delta_{1,n}^{+}(k)\}_{n\geq 1},$
and a return time function by
$\tau_{k}(x)=\inf\{n\geq 1\>:\>T_{k,k+n-1}(x)\in Y_{k+n}\}.$
Since $\tau_{k}$ is constant on the elements of ${\mathcal{P}}_{k}$, (NU:2) holds. Using the convexity of the two branches together with properties of the first derivatives in (3.3), we see that $|F_{a}(y)-F_{a}(y^{\prime})|\geq\Lambda|y-y^{\prime}|>0$ holds for some $\Lambda>1$ whenever $y,y^{\prime}\in a$ and $a\in{\mathcal{P}}_{k}$.
Proposition 3.16.
There exists $K>0$ such that for any $k\geq 1$ and $a\in{\mathcal{P}}_{k}$,
$\zeta=\frac{d(F_{a})_{*}(m_{k}|_{a})}{dm_{k+\tau_{k}(a)}}\quad\text{satisfies}\quad|\zeta|_{{\mathrm{LL}},k+\tau_{k}(a)}\leq K.$
Let $\mu$ be a probability measure on $X$ with a positive Hölder continuous density. Then, for sufficiently large $K_{1}$ in the definition of regularity, $\mu$ is regular w.r.t. $T_{k},T_{k+1},\ldots$ for all $k\geq 1$.
Remark 3.17.
The case of nonnegative Hölder densities can be treated by using the decomposition (3.13).
Proof of Proposition 3.16.
The estimate on $|\zeta|_{{\mathrm{LL}},k+\tau_{k}(a)}$ follows by the distortion bounds in [20, Lemma 3.10], using a similar argument as in the proof of Proposition 3.11. The regularity of $\mu$ can be shown as in the proof of Proposition 3.13. ∎
Hence, (NU:3) holds. Again (NU:4) is clear since all maps in (3.14) are expanding. For (NU:5), we observe that, for some constant $\delta_{\#}>0$ independent of $k$,
$m_{k}(\tau_{k}=2)\geq|\delta_{2}^{-}(k)|\geq\delta_{\#}\quad\text{and}\quad m_{k}(\tau_{k}=3)\geq|\delta_{3}^{-}(k)|\geq\delta_{\#}.$
[…]
Proposition 3.18.
We have the following uniform tail bounds:
$\sup_{k}m_{k}(\tau_{k}\geq n)=O\bigl{(}n^{-\frac{1}{\eta_{+}(\gamma_{+}-1)}}\bigr{)}\quad\text{and}\quad\lambda(\tau_{k}\geq n)=O(n^{-\frac{1}{\gamma_{+}-1}}).$
Proof.
If $\mu$ is a measure on $X$, then for $n\geq 2$,
$\mu(\tau_{k}\geq n)$ $=\mu\biggl{(}\bigcup_{\ell\geq n}\Delta^{\pm}_{\ell}(k)\biggr{)}+\mu\biggl{(}\bigcup_{\ell\geq n}\delta^{\pm}_{\ell}(k)\biggr{)}+\mu\biggl{(}\bigcup_{\ell\geq n-1}\delta^{\pm}_{1,\ell}(k)\biggr{)}=I+II+III.$ (3.18)
First, suppose that $\mu=m_{k}$. Then, we have $I=0$. By [20, Lemma 3.3], $II=O(n^{-1/\eta_{+}(\gamma_{+}-1)})$. Moreover,
$\lambda\biggl{(}\bigcup_{\ell\geq n-1}\delta^{+}_{1,\ell}(k)\biggr{)}\leq\lambda\biggl{(}T_{k}^{-1}\bigcup_{\ell\geq n-1}\delta^{+}_{\ell}(k+1)\biggr{)}\leq C\lambda\biggl{(}\bigcup_{\ell\geq n-1}\delta^{+}_{\ell}(k+1)\biggr{)}\leq C^{\prime}n^{-1/\eta_{+}(\gamma_{+}-1)},$
where $C^{\prime}$ is independent of $k$. The same estimate holds for $\delta^{-}_{1,\ell}(k)$ in place of $\delta^{+}_{1,\ell}(k)$, so that $III=O(n^{-1/\eta_{+}(\gamma-1)})$. This gives the desired tail bound for $\mu=m_{k}$.
Finally, suppose that $\mu=\lambda$. By [20, Sublemma 3.2 and equation (53)], $I=O(n^{-1/(\gamma_{+}-1)})$, so the tail bound for $\lambda$ follows. ∎
To complete the proof of Theorem 3.14, it remains to show the following.
Proposition 3.19.
Assume (3.17). Then, for each $k\geq 1$, $h^{k}$ in (2.1) satisfies
$h^{k}(\ell)=O\bigl{(}(1\vee(\ell-\Theta k))^{-1/(\gamma-1)}\bigr{)},$
where the implied constant is independent of $k$, and $\Theta=c_{0}\kappa(1-\kappa)^{-1}$ for some absolute constant $c_{0}>0$. Further, for any probability measure $\mu$ on $X$ with bounded density, $\mu(\tau_{1}\geq n)=O(n^{-1/(\gamma-1)})$.
Proof.
Let $(T_{k})$ be a sequence of maps as in Theorem 3.14. As in the proof of Theorem 1.1, we let $\widetilde{T}_{j}=T_{n_{j}}$ where $(T_{n_{j}})$ is the sub-sequence of maps whose parameters satisfy $\gamma_{n_{j}}\leq\gamma$. Set
$G_{\ell}(k)$ =#($=\#\bigl{(}\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr{)},$
$\widetilde{\tau}_{k}(\ell)$ $=\inf\{\ell\geq 1\>:\>\widetilde{T}_{m(k),m(k)+\ell-1}(x)\in Y_{n_{m(k)+\ell}}\},$
$m(k)$ $=\min\{j\geq 1\>:\>n_{j}\geq k\}.$
By the definition of $\Delta_{n}^{-}(k)$ we have $\cup_{j\geq\ell}\Delta_{j}^{-}(k)\subset\{\widetilde{\tau}_{k}\geq G_{\ell}(k)\}$. Consequently, using the inclusions $\Delta_{\ell}^{+}(k)\subset T_{k}^{-1}\Delta_{\ell-1}^{-}(k+1)$, $\delta^{\pm}_{\ell}(k)\subset T_{k}^{-1}\Delta_{\ell-1}^{+}(k+1)$, and $\delta_{1,\ell}^{\pm}(k)\subset T_{k}^{-1}\delta_{\ell}^{+}(k+1)$ together with the assumption that the density of $\mu$ is bounded, following the proof of Proposition 3.18, we obtain
$\mu(\tau_{k}\geq\ell)=I+II+III\leq C\sum_{j=0}^{3}G_{\ell-j}(k+j)^{-1/(\gamma-1)}\quad\forall\ell\geq 3,$
where $C$ is independent of $k$. Finally, as in (3.1), by invoking (3.17) we see that for some absolute constant $c_{0}$,
$G_{\ell}(k)\geq a(1-\kappa)\biggl{(}\ell-\frac{c_{0}\kappa}{1-\kappa}k\biggr{)},$
whenever $k\geq N$ and $\ell$ is sufficiently large (depending on $\kappa$). ∎
4. Proof of Theorem 2.6
Define $C_{h}=2e^{K_{2}\operatorname{diam}X}$ and
$h_{n}^{k}(\ell)=C_{h}\bigl{(}h^{k}(n+\ell)+h^{k+1}(n+\ell-1)+\ldots+h^{k+n}(\ell)\bigr{)}.$ (4.1)
Proposition 4.1.
Let $k\geq 1$. Suppose that $\mu$ is a probability measure on $Y_{k}$ with $|\mu|_{{\mathrm{LL}},k}\leq K_{2}$.
  1. (a)
    For every $n\geq 0$, the measure $(T_{k,k+n-1})_{*}\mu$ is regular with tail bound $\frac{1}{2}h^{k}_{n}$ with respect to $T_{k+n},T_{k+n+1},\ldots$
  2. (b)
    There is a constant $\theta\in(0,1)$, depending only on $K_{1}$, $K_{2}$, $\operatorname{diam}X$ and $\delta_{0}$, such that for every $n\geq n_{0}$,
    $(T_{k,k+n-1})_{*}\mu=\theta m_{k+n}+(1-\theta)\mu^{\prime},$
    where $\mu^{\prime}$ is a regular probability measure with tail bound $h^{k}_{n}$ with respect to $T_{k+n},T_{k+n+1},\ldots$
Proof.
We prove (a) first. Fix $n\geq 0$ and $k\geq 1$. For each $0\leq j\leq n$, define
$Y^{\prime}_{j}$ =$=\{y\in Y_{k+n-j}:T_{k+n-j,k+\ell-1}(y)\notin Y_{k+\ell}\text{ for all $n-j<\ell\leq n$}\},$
$Y_{j}^{\prime\prime}$ $=Y_{k}\cap T_{k,k+n-j-1}^{-1}(Y_{j}^{\prime})$
=$=\{y\in Y_{k}:T_{k,k+n-j-1}(y)\in Y_{k+n-j}\text{ and $T_{k,k+\ell-1}(y)\notin Y_{k+\ell}$ for all $n-j<\ell\leq n$}\}.$
In particular, $Y_{0}^{\prime}=Y_{k+n}$ and $Y^{\prime\prime}_{0}=\{y\in Y_{k}:T_{k,k+n-1}(y)\in Y_{k+n}\}$.
Observe that the sets $Y_{j}^{\prime\prime}$ form a partition of $Y_{k}$, so we can write
$(T_{k,k+n-1})_{*}\mu=\sum_{j=0}^{n}(T_{k,k+n-1})_{*}\mu_{j},$ (4.2)
where $\mu_{j}$ is the restriction of $\mu$ to $Y_{j}^{\prime\prime}$.
Next, define $\nu_{j}=((T_{k,k+n-j-1})_{*}\mu)|_{Y_{k+n-j}}$ for $0\leq j\leq n$, and note that for all $B\subset X$,
$(T_{k+n-j,k+n-1})_{*}(\nu_{j}|_{Y_{j}^{\prime}})(B)$ $=\nu_{j}(Y_{j}^{\prime}\cap T_{k+n-j,k+n-1}^{-1}(B))$
$=\mu(T_{k,k+n-j-1}^{-1}(Y_{j}^{\prime})\cap T_{k,k+n-1}^{-1}(B))$
$=(T_{k,k+n-1})_{*}\mu_{j}(B).$
Hence, $(T_{k+n-j,k+n-1})_{*}(\nu_{j}|_{Y_{j}^{\prime}})=(T_{k,k+n-1})_{*}\mu_{j}$.
By Proposition 2.5, $|\nu_{j}|_{{\mathrm{LL}},k+n-j}\leq K_{2}$. Since $\nu_{j}(X)\leq 1$, we deduce that $\nu_{j}\leq e^{K_{2}\operatorname{diam}X}m_{k+n-j}$ and thus the tail of $\nu_{j}$ with respect to $T_{k+n-j},T_{k+n-j+1},\ldots$ is bounded by $e^{K_{2}\operatorname{diam}X}h^{k+n-j}$. Observe that $(T_{k+n-j,k+n-1})_{*}(\nu_{j}|_{Y_{j}^{\prime}})$ inherits the tail bound from $\nu_{j}$ with a time shift, namely $(T_{k+n-j,k+n-1})_{*}(\nu_{j}|_{Y_{j}^{\prime}})$ has tail bound $e^{K_{2}\operatorname{diam}X}h^{k+n-j}(\cdot+j)$ with respect to $T_{k+n},T_{k+n+1},\ldots$ By (4.2), it follows that $(T_{k,k+n-1})_{*}\mu$ has tail bound $e^{K_{2}\operatorname{diam}X}\sum_{j=0}^{n}h^{k+n-j}(\cdot+j)$ with respect to $T_{k+n},T_{k+n+1},\ldots$, as required. Moreover, $(T_{k,k+n-1})_{*}\mu$ is regular w.r.t. $T_{k+n},T_{k+n+1},\ldots$ by Proposition 2.5.
It remains to prove (b). Let $\theta_{0}\in(0,1)$ be such that for every $\theta^{\prime}\in[0,\theta_{0}]$, every measure $\rho$ on $Y_{k}$ with $|\rho|_{{\mathrm{LL}},k}\leq K_{1}$ can be written as $\rho=\rho(Y_{k})\theta^{\prime}m_{k}+\rho^{\prime}_{k}$ with $|\rho^{\prime}_{k}|_{{\mathrm{LL}},k}\leq K_{2}$. Such $\theta_{0}$ exists and only depends on $K_{1},K_{2}$ and $\operatorname{diam}(X)$, see [15, Lemma 3.4].
Suppose that $n\geq n_{0}$, and let $\rho_{n}=\bigl{(}(T_{k,k+n-1})_{*}\mu\bigr{)}\big{|}_{Y_{k+n}}$. By Proposition 2.5, $\bigl{|}\rho_{n}\bigr{|}_{{\mathrm{LL}},k+n}\leq K_{2}$, and by (NU:5), $\rho_{n}(Y_{k+n})\geq\delta_{0}e^{-\operatorname{diam}(X)K_{2}}$. Let $\theta=\min\{\theta_{0}\delta_{0}e^{-\operatorname{diam}(X)K_{2}},1/2\}$. Then
$\rho_{n}=\theta m_{k+n}+\rho^{\prime}_{k+n}\qquad\text{with }|\rho^{\prime}_{k+n}|_{{\mathrm{LL}},k+n}\leq K_{2}.$
Define
$\mu^{\prime}=(1-\theta)^{-1}\bigl{(}(T_{k,k+n-1})_{*}\mu-\theta m_{k+n}\bigr{)}=(1-\theta)^{-1}\Bigl{(}\rho^{\prime}_{k+n}+\bigl{(}(T_{k,k+n-1})_{*}\mu\bigr{)}\big{|}_{X\setminus Y}\Bigr{)}.$
Then $\mu^{\prime}$ is a probability measure and $(T_{k,k+n-1})_{*}\mu=\theta m_{k+n}+(1-\theta)\mu^{\prime}$. By Proposition 2.5, $\rho^{\prime}_{k+n}$ and $\bigl{(}(T_{k,k+n-1})_{*}\mu\bigr{)}\big{|}_{X\setminus Y}$ are regular measures w.r.t. $T_{k+n},T_{k+n+1},\ldots$ hence so is $\mu^{\prime}$. To bound the tail of $\mu^{\prime}$, we note that $\mu^{\prime}\leq(1-\theta)^{-1}(T_{k,k+n-1})_{*}\mu$ with $(1-\theta)^{-1}\leq 2$ and apply the bound on the tail of $(T_{k,k+n-1})_{*}\mu$ from (a). ∎
Fix $\theta$ as in Proposition 4.1. We can translate the decomposition of a pushforward $(T_{k,k+N-1})_{*}\mu=\theta m_{k+N}+(1-\theta)\mu^{\prime}$ into a corresponding decomposition of $\mu$ by simply pulling back measures:
Corollary 4.2.
Suppose that $N\geq 0$, $k\geq 1$, and $\mu$ is a probability measure on $X$ such that $(T_{k,k+N-1})_{*}\mu$ is supported on $Y_{k+N}$ and $|(T_{k,k+N-1})_{*}\mu|_{{\mathrm{LL}},k+N}\leq K_{2}$. Then for every $n\geq N+n_{0}$,
$\mu=\theta\mu_{n}+(1-\theta)\mu^{\prime}_{n},$
where $\mu_{n}$, $\mu^{\prime}_{n}$ are probability measures with $(T_{k,k+n-1})_{*}\mu_{n}=m_{k+n}$ and is $(T_{k,k+n-1})_{*}\mu^{\prime}_{n}$ regular with tail bound $h_{n-N}^{k+N}$ w.r.t. $T_{k+n},T_{k+n+1},\ldots$
Proof.
Fix $n\geq N+n_{0}$. Proposition 4.1 gives the decomposition $(T_{k,k+n-1})_{*}\mu=(T_{k+N,k+n-1})_{*}((T_{k,k+N-1})_{*}\mu)=\theta m_{k+n}+(1-\theta)\mu^{\prime}$, where $\mu^{\prime}$ is a regular probability measure with tail bound $h_{n-N}^{k+N}$ w.r.t. $T_{k+n},T_{k+n+1},\ldots$ Define $\mu_{n}$ and $\mu_{n}^{\prime}$ by
$d\mu_{n}=\Bigl{(}\frac{dm_{k+n}}{d(T_{k,k+n-1})_{*}\mu}\circ T_{k,k+n-1}\Bigr{)}\,d\mu\quad\text{and}\quad d\mu^{\prime}_{n}=\Bigl{(}\frac{d\mu^{\prime}}{d(T_{k,k+n-1})_{*}\mu}\circ T_{k,k+n-1}\Bigr{)}\,d\mu.$
It is straightforward that $\mu=\theta\mu_{n}+(1-\theta)\mu^{\prime}_{n}$ with $(T_{k,k+n-1})_{*}\mu_{n}=m_{k+n}$ and $(T_{k,k+n-1})_{*}\mu^{\prime}_{n}=\mu^{\prime}$. This is the desired decomposition. ∎
Lemma 4.3.
Suppose that $\mu$ is a regular probability measure with tail bound $r$ w.r.t. $T_{k},T_{k+1},\ldots$ where $k\geq 1$. Suppose that $r(n)$ is nonincreasing, $r(1)=1$ and $\lim_{n\to\infty}r(n)=0$. Then
$\mu=\sum_{j=n_{0}+1}^{\infty}\alpha_{j}[\theta\mu_{j}+(1-\theta)\mu^{\prime}_{j}],$
where $\alpha_{j}=r(j-n_{0})-r(j+1-n_{0})$ and $\mu_{j}$, $\mu^{\prime}_{j}$ are probability measures such that $(T_{k,k+j-1})_{*}\mu_{j}=m_{k+j}$ and $(T_{k,k+j-1})_{*}\mu^{\prime}_{j}$ is regular with tail bound $h_{j}^{k}$ w.r.t. $T_{k+j},T_{k+j+1},\ldots$
Proof.
Recall the partition ${\mathcal{P}}_{k}$ of $X$ corresponding to $T_{k},T_{k+1},\ldots$ For all $n\geq 1$, define
$a_{k,n}=\cup\{a\in{\mathcal{P}}_{k}:\tau_{k}(a)=n\}.$
Let $\nu_{n}=\mu|_{a_{k,n}}$. Then, for each $n\geq 1$, the measure $(T_{k,k+n-1})_{*}\nu_{n}$ is supported on $Y_{k+n}$ and satisfies $|(T_{k,k+n-1})_{*}\nu_{n}|_{{\mathrm{LL}},k+n}\leq K_{1}$. By Corollary 4.2, for each $\ell\geq n+n_{0}$,
$\nu_{n}=\mu(a_{k,n})\bigl{[}\theta\nu_{n,\ell}+(1-\theta)\nu_{n,\ell}^{\prime}\bigr{]},$ (4.3)
where $\nu_{n,\ell}$, $\nu_{n,\ell}^{\prime}$ are probability measures with $(T_{k,k+\ell-1})_{*}\nu_{n,\ell}=m_{k+\ell}$ and $(T_{k,k+\ell-1})_{*}\nu_{n,\ell}^{\prime}$ is regular with tail bound $h_{\ell-n}^{k+n}$ w.r.t. $T_{k+\ell},T_{k+\ell+1},\ldots$
We observe that
$\sum_{n=1}^{\ell}\mu(a_{k,n})\geq r(1)-r(\ell+1)\quad\text{ and }\quad\sum_{n=1}^{\infty}\mu(a_{k,n})=r(1)-\lim_{\ell\to\infty}r(\ell)=1.$
By [15, Proposition 4.7] there exist nonnegative constants $\xi_{\ell,n}$, $1\leq n\leq\ell<\infty$, such that
[…]
For $j\geq n_{0}+1$, define $\chi_{j}=\sum_{n=1}^{j-n_{0}}\xi_{j-n_{0},n}\nu_{n}$. Then $\mu=\sum_{j=n_{0}+1}^{\infty}\chi_{j}$ and $\chi_{j}(X)=\alpha_{j}$. From (4.3), we have
$\chi_{j}=\alpha_{j}[\theta\mu_{j}+(1-\theta)\mu^{\prime}_{j}]$
with $\mu_{j}=\alpha_{j}^{-1}\sum_{n=1}^{j-n_{0}}\xi_{j-n_{0},n}\mu(a_{k,n})\nu_{n,j}$ and $\mu^{\prime}_{j}=\alpha_{j}^{-1}\sum_{n=1}^{j-n_{0}}\xi_{j-n_{0},n}\mu(a_{k,n})\nu^{\prime}_{n,j}$. It is possible that $\alpha_{j}=0$, but this does not create problems and we ignore it for simplicity.
It remains to observe that $\mu_{j}$ and $\mu^{\prime}_{j}$ are probability measures with $(T_{k,k+j-1})_{*}\mu_{j}=m_{k+j}$ and $(T_{k,k+j-1})_{*}\mu^{\prime}_{j}$ is regular with tail bound $\max\{h^{k+1}_{j-1},\ldots,h^{k+j-n_{0}}_{n_{0}}\}\leq h_{j-1}^{k+1}\leq h_{j}^{k}$ w.r.t. $T_{k+j},T_{k+j+2},\ldots$
Similar to Corollary 4.2, we obtain the following result:
Corollary 4.4.
Suppose that $N\geq 0$, $k\geq 1$, and $\mu$ is a probability measure such that $(T_{k,k+N-1})_{*}\mu$ is regular and has tail bound $r$ w.r.t. $T_{k+N},T_{k+N+1},\ldots$, where $r(n)$ is nonincreasing, $r(1)=1$ and $\lim_{n\to\infty}r(n)=0$. Then
$\mu=\sum_{j=n_{0}+1}^{\infty}\alpha_{j}[\theta\mu_{j}+(1-\theta)\mu^{\prime}_{j}],$ (4.4)
where $\alpha_{j}=r(j-n_{0})-r(j+1-n_{0})$, and $\mu_{j}$, $\mu^{\prime}_{j}$ are probability measures such that $(T_{k,k+N+j-1})_{*}\mu_{j}=m_{k+N+j}$ and $(T_{k,k+N+j-1})_{*}\mu^{\prime}_{j}$ is regular with tail bound $h^{k+N}_{j}$ w.r.t. $T_{k+N+j},T_{k+N+j+1},\ldots$
Further, we suppose that $r$ is nonnegative with $\lim_{n\to\infty}r(n)=0$ and define
${\hat{r}}(n)=\min\{1,r(1),\ldots,r(n)\}.$
This way, ${\hat{r}}$ is nonincreasing and ${\hat{r}}(1)=1$; for a probability measure, the tail bounds $r$ and ${\hat{r}}$ are equivalent. Similarly define ${\hat{h}}^{k}_{n}$.
Let $X_{1},X_{2},\ldots$ be random variables with values in $\{n_{0},n_{0}+1,\ldots\}$, such that for all $\ell\geq n_{0}$,
$\operatorname{\mathbb{P}}(X_{1}\geq\ell)$ $={\hat{r}}(\ell-n_{0}),$ (4.5)
$\operatorname{\mathbb{P}}(X_{j+1}\geq\ell\mid X_{1},\ldots,X_{j})$ $={\hat{h}}^{k+X_{1}+\ldots+X_{j-1}}_{X_{j}}(\ell-n_{0})\quad\text{ for }j\geq 1.$
Let $\tau$ be a geometrically distributed random variable with parameter $\theta$, independent of $\{X_{j}\}$. That is,
$\operatorname{\mathbb{P}}(\tau=\ell)=(1-\theta)^{\ell-1}\theta\quad\forall\ell\in\{1,2,\ldots\}.$
Define
$S=X_{1}+\ldots+X_{\tau}.$
Note that $S$ depends on $k$.
Lemma 4.5.
Suppose that $\mu$ is a regular probability measure with tail bound $r$ w.r.t. $T_{k},T_{k+1},\ldots$ Then there exists a decomposition
$\mu=\sum_{n=1}^{\infty}\operatorname{\mathbb{P}}(S=n)\mu_{n},$
where $\mu_{n}=\mu_{n}^{(k)}$ are probability measures such that $(T_{k,k+n-1})_{*}\mu_{n}=m_{k+n+1}$.
Proof.
By iterating the decomposition in (4.4), we can express $\mu$ as an infinite series
$\mu$ $=\theta\sum_{i>n_{0}}\alpha_{i}\mu_{i}+(1-\theta)\theta\sum_{i,j>n_{0}}\alpha_{i,j}\mu_{i,j}+(1-\theta)^{2}\theta\sum_{i,j,\ell>n_{0}}\alpha_{i,j,\ell}\mu_{i,j,\ell}+\cdots$ (4.6)
as follows.
  • Step 1: Apply (4.4) to $\mu$:
    $\mu=\sum_{i>n_{0}}\alpha_{i}[\theta\mu_{i}+(1-\theta)\mu_{i}^{\prime},]$
    where $\alpha_{i}=\hat{r}(i-n_{0})-\hat{r}(i+1-n_{0})$, $(T_{k,k+i-1})_{*}\mu_{i}=m_{k+i}$, and $(T_{k,k+i-1})_{*}\mu_{i}^{\prime}$ is regular with tail bound $\hat{h}^{k}_{i}$ w.r.t. $T_{k+i},T_{k+i+1},\ldots$
  • Step 2: Apply (4.4) to each $\mu_{i}^{\prime}$:
    $\mu=\theta\sum_{i>n_{0}}\alpha_{i}\mu_{i}+(1-\theta)\theta\sum_{i,j>n_{0}}\alpha_{i,j}\mu_{i,j}+(1-\theta)^{2}\sum_{i,j>n_{0}}\alpha_{i,j}\mu_{i,j}^{\prime},$
    where $\alpha_{i,j}=\alpha_{i}(\hat{h}^{k}_{i}(j-n_{0})-\hat{h}^{k}_{i}(j+1-n_{0}))$, $(T_{k,k+i+j-1})_{*}\mu_{i,j}=m_{k+i+j}$, and $(T_{k,k+i+j-1})_{*}\mu_{i,j}^{\prime}$ is regular with tail bound $h_{j}^{k+i}$ w.r.t. $T_{k+i+j},T_{k+i+j+1},\ldots$
  • Step 3: Apply (4.4) to each $\mu_{i,j}^{\prime}$:
    $\mu$ $=\theta\sum_{i>n_{0}}\alpha_{i}\mu_{i}+(1-\theta)\theta\sum_{i,j>n_{0}}\alpha_{i,j}\mu_{i,j}+(1-\theta)^{2}\theta\sum_{i,j,\ell>n_{0}}\alpha_{i,j,\ell}\mu_{i,j,\ell}$
    $+(1-\theta)^{3}\sum_{i,j,\ell>n_{0}}\alpha_{i,j,\ell}\mu_{i,j,\ell}^{\prime},$
    where $\alpha_{i,j,\ell}=\alpha_{i,j}(\hat{h}^{k+j}_{i}(\ell-n_{0})-\hat{h}^{k+j}_{i}(\ell+1-n_{0}))$, $(T_{k,k+i+j+\ell-1})_{*}\mu_{i,j,\ell}=m_{k+i+j+\ell}$, and $(T_{k,k+i+j+\ell-1})_{*}\mu_{i,j,\ell}^{\prime}$ is regular with tail bound $h_{\ell}^{k+i+j}$ w.r.t. $T_{k+i+j+\ell},T_{k+i+j+\ell+1},\ldots$
  • And so on…
By induction, we find that for each $n\geq 1$ and $j_{1},j_{2},\ldots,j_{n}\geq n_{0}$,
$(1-\theta)^{n-1}\theta\alpha_{j_{1},j_{2},\ldots,j_{n}}=\operatorname{\mathbb{P}}(\tau=n,X_{1}=j_{1},\ldots,X_{n}=j_{n}).$
Grouping the terms in (4.6) by the sum of indices, we obtain the required decomposition with
$\mu_{n}=\sum_{\begin{subarray}{c}k\geq 1\\ j_{1}+\cdots+j_{k}=n\end{subarray}}(1-\theta)^{k-1}\theta\alpha_{j_{1},\ldots,j_{k}}\mu^{\prime}_{j_{1},\ldots,j_{k}}\Bigg{/}\sum_{\begin{subarray}{c}k\geq 1\\ j_{1}+\cdots+j_{k}=n\end{subarray}}(1-\theta)^{k-1}\theta\alpha_{j_{1},\ldots,j_{k}}.$
To complete the proof of Theorem 2.6, it remains to estimate the tails $\operatorname{\mathbb{P}}(S\geq n)$, as is done in the following proposition.
Proposition 4.6.
There exists $\Theta\in(0,1)$ such that the following holds for any sequence of numbers $\Theta_{j}\in[0,1)$ with $\sup_{j\geq 1}\Theta_{j}\leq\Theta$. Suppose that $\mu$ is a regular probability measure with tail bound $r$ w.r.t. $T_{k},T_{k+1},\ldots$ for $k\geq 1$, and that for some $0<\beta^{\prime}\leq\beta$, $\beta>1$, $C_{\beta},C_{\beta}^{\prime}\geq 1$,
$h^{j}(n)\leq C_{\beta}(1\vee(n-\Theta_{j}j))^{-\beta}\quad\text{and}\quad r(n)\leq C_{\beta}^{\prime}(1\vee(n-\Theta_{k}k))^{-\beta^{\prime}}$
hold for all $j\geq k$ and $n\geq 1$. Then, for all $n\geq 1$,
$\operatorname{\mathbb{P}}(S\geq n)\leq C_{\beta}^{\prime}C(\Theta_{k}^{*}k+1)^{\beta^{\prime}}n^{-\beta^{\prime}},$
where $\Theta_{k}^{*}=\sup_{j\geq k}\Theta_{j}$, and $C>0$ is a constant depending only on $n_{0},\theta,\Theta,C_{h},C_{\beta},\beta,\beta^{\prime}$.
Proof.
Suppose, without loss of generality, that $h^{j}$ are nonincreasing, so that ${\hat{h}}^{j}_{n}(\ell)\leq h^{j}_{n}(\ell)$.
Denote $S_{j}=X_{1}+\cdots+X_{j}$; then $S=S_{\tau}$. We regard $n\geq 1$ and $j\geq 1$ as fixed and write
$\operatorname{\mathbb{P}}(S_{j+1}\geq n)=\sum_{\ell=1}^{n+1}H_{\ell},$ (4.7)
where
$H_{\ell}$ $=\operatorname{\mathbb{P}}(X_{j+1}\geq n-\ell\mid S_{j}=\ell)\operatorname{\mathbb{P}}(S_{j}=\ell),\quad 1\leq\ell\leq n,$
$H_{n+1}$ $=\operatorname{\mathbb{P}}(S_{j}>n).$
From (4.1),
$h^{k+S_{j-1}}_{X_{j}}\leq h^{k}_{S_{j}}.$
Using this inequality together with (4.5),
$\operatorname{\mathbb{P}}(X_{j+1}\geq n-\ell\mid S_{j}=\ell)\leq h^{k}_{\ell}(n-\ell-n_{0})=C_{h}\sum_{i=0}^{\ell}h^{k+i}(n-i-n_{0}),$
where we use the convention $h^{i}(\ell)=0$ for $\ell\leq 0$.
To estimate (4.7), we decompose the sum $\sum_{1\leq\ell\leq n+1}H_{\ell}$ intro three parts. To this end, fix $b\in(1/2,1)$ by $b=\bigl{(}1+(1-\theta)^{\frac{1}{2\beta^{\prime}}}\bigr{)}/2$ and let $A\geq 1$ be a large integer whose value will be specified later. Restrict to $\Theta$ sufficiently small so that $01-b-\Theta b>$ 0. Let $C$ denote various constants which depend only on $A,n_{0},\theta,\Theta,C_{h},C_{\beta},\beta,\beta^{\prime}$. In particular, $C$ is independent of $k$.
Define
$R_{j}=\sup_{n\geq 1}n^{\beta^{\prime}}\operatorname{\mathbb{P}}(S_{j}\geq n).$
Suppose that
$n>\max\biggl{\{}\frac{8(A+n_{0})}{1-b-\Theta b},b^{-1}\frac{2\Theta_{k}^{*}k}{1-b-\Theta b}\biggr{\}}.$ (4.8)
By our assumption on $h^{j}$,
$\sum_{\ell\leq A}H_{\ell}$ $\leq\sup_{\ell\leq A}\operatorname{\mathbb{P}}(X_{k+1}\geq n-\ell\mid S_{j}=\ell)\leq C_{h}\sum_{i\leq A}h^{k+i}(n-i-n_{0})$ (4.9)
$\leq C_{h}C_{\beta}A(n-2A-n_{0}-\Theta_{k}^{*}k)^{-\beta}$
$\leq Cn^{-\beta}.$
In particular,
$\sum_{i\leq A}h^{k+i}(n-i-n_{0})\leq Cn^{-\beta}.$ (4.10)
Next,
$\sum_{bn\leq\ell\leq n+1}H_{\ell}\leq\operatorname{\mathbb{P}}(S_{j}\geq bn)\leq R_{j}b^{-\beta^{\prime}}n^{-\beta^{\prime}}.$ (4.11)
Further, it is a direct verification that
$\sum_{A<\ell\leq bn}h^{k+\ell}(n-\ell-n_{0})\leq C\sum_{A<\ell\leq bn}(n-\ell-n_{0}-\Theta_{k}^{*}(k+\ell))^{-\beta}\ell^{-\beta^{\prime}}\leq A_{n}n^{-\beta^{\prime}},$ (4.12)
where $\lim_{A\to\infty,\,n\to\infty}A_{n}=0$. Now we make the promised choice of $A$, fixing it sufficiently large so that
$b^{-\beta^{\prime}}+C_{h}A_{n}\leq(1-\theta)^{-1/2}$ (4.13)
for all $n$ satisfying (4.8).
Using summation by parts together with (4.10) and (4.12),
$\sum_{A\leq\ell<bn}H_{\ell}$ $\leq\sum_{A\leq\ell<bn}C_{h}\sum_{i\leq\ell}h^{k+i}(n-i-n_{0})\operatorname{\mathbb{P}}(S_{j}=\ell)$ (4.14)
$=-C_{h}\operatorname{\mathbb{P}}(S_{j}\geq bn)\sum_{i\leq bn}h^{k+i}(n-i-n_{0})$
$\quad\,+C_{h}\operatorname{\mathbb{P}}(S_{j}\geq A)\sum_{i\leq A}h^{k+i}(n-i-n_{0})$
$\quad\,+C_{h}\sum_{A<\ell\leq bn}h^{k+\ell}(n-\ell-n_{0})\operatorname{\mathbb{P}}(S_{j}\geq\ell)$
$\leq Cn^{-\beta}+C_{h}A_{n}R_{j}n^{-\beta^{\prime}}.$
Assembling (4.7), (4.9), (4.11) and (4.14), we obtain
$\operatorname{\mathbb{P}}(S_{j+1}\geq n)$ $\leq Cn^{-\beta}+R_{j}n^{-\beta^{\prime}}(b^{-\beta^{\prime}}+C_{h}A_{n})$
$\leq Cn^{-\beta}+R_{j}n^{-\beta^{\prime}}(1-\theta)^{-1/2}$
for all $n$ satisfying (4.8), in particular for $n\geq C\Theta_{k}^{*}k$. Hence for all $n\geq 1$,
$\operatorname{\mathbb{P}}(S_{j+1}\geq n)\leq Cn^{-\beta}+R_{j}n^{-\beta^{\prime}}(1-\theta)^{-1/2}+C(\Theta_{k}^{*}k)^{\beta^{\prime}}n^{-\beta^{\prime}}.$
Then
$R_{j+1}\leq C+C(\Theta_{k}^{*}k)^{\beta^{\prime}}+(1-\theta)^{-1/2}R_{j}.$
By induction,
$R_{j}\leq C_{\beta}^{\prime}C(\Theta_{k}^{*}k+1)^{\beta^{\prime}}(1-\theta)^{-j/2}.$
This extends to
$\operatorname{\mathbb{P}}(S_{j}\geq n)\leq C_{\beta}^{\prime}C_{2}(\Theta_{k}^{*}k+1)^{\beta^{\prime}}(1-\theta)^{-j/2}n^{-\beta^{\prime}}.$
Since $\tau$ and $S_{j}$ are independent and $\operatorname{\mathbb{P}}(\tau=j)=(1-\theta)^{j-1}\theta$,
$\operatorname{\mathbb{P}}(S\geq n)$ $=\sum_{j\geq 1}\operatorname{\mathbb{P}}(S_{j}\geq n)\operatorname{\mathbb{P}}(\tau=j)$
[…]
as desired. ∎
Appendix A Proof of Proposition 2.1
Assume that conditions (i) and (ii) in Proposition 2.1 hold. Let $N_{1}=\max\{p_{1}^{2},\ldots,p_{M}^{2}\}$. By [25], for all $n\geq N_{1}$ we can write $n=\sum_{j=1}^{M}a_{j}p_{j}$ for some integers $a_{j}\geq 0$. For $N_{1}\leq N<n$, we have
$(T_{k,k+n-1})_{*}m_{k}=(T_{k+n-N,k+n-1})_{*}\nu,$
where $\nu=(T_{k,k+n-N-1})_{*}m_{k}$. By Proposition 2.5, $\nu$ is regular w.r.t $T_{k+n-N},T_{k+n-N+1},\ldots$, and $\bigl{|}\nu|_{Y_{k+n-N}}\bigr{|}_{{\mathrm{LL}},k+n-N}\leq K_{1}$.
Let $N=N_{1}+N_{\#}$. Condition (i) in Proposition 2.1 reads
$\sup_{k,j\geq 1}(T_{k,k+j-1})_{*}m_{k}(\tau_{k+j}\geq N_{\#})\leq 1/2.$ (A.1)
Using the regularity of $\nu$, we obtain
$(T_{k,k+n-1})_{*}m_{k}(Y_{k+n})$
[…]
$=\sum_{\ell\leq N_{\#}}(T_{k+n-N+\ell,k+n-1})_{*}(T_{k+n-N,k+n-N+\ell-1})_{*}(\nu|_{\{\tau_{k+n-N}=\ell\}})(Y_{k+n})$
$\geq\sum_{\ell\leq N_{\#}}e^{-\text{diam}(X)K_{1}}(T_{k+n-N,k+n-N+\ell-1})_{*}(\nu|_{\{\tau_{k+n-N}=\ell\}})(Y_{k+n-N+\ell})$
[…]
As in [15, Proposition 4.4], using our assumption $m_{k}(\tau_{k}=p_{m})\geq\delta_{\#}$ together with Proposition 2.5, we find that for any $\ell\leq N_{\#}$,
$(T_{k+n-N+\ell,k+n-1})_{*}m_{k+n-N+\ell}(Y_{k+n})\geq(\delta_{\#}e^{-\operatorname{diam}(X)K_{1}})^{N}.$
Summing over $\ell$, we arrive at the lower bound
$(T_{k,k+n-1})_{*}m_{k}(Y_{k+n})\geq(\delta_{\#}e^{-\operatorname{diam}(X)K_{1}})^{N+1}\nu(\tau_{k+n-N}\leq N_{\#}).$
By (A.1), $\nu(\tau_{k+n-N}\leq N_{\#})\geq 1/2$. Therefore,
$(T_{k,k+n-1})_{*}m_{k}(Y_{k+n})\geq(\delta_{\#}e^{-\operatorname{diam}(X)K_{1}})^{N+1}/2,$
whenever $n>N_{1}+N_{\#}$. Thus, condition (NU:5) holds.
[…]
$(T_{k,k+j-1})_{*}m_{k}(\tau_{k+j}>N_{\#})$ $\leq\frac{1}{2}h_{j}^{k}(N_{\#}+1)$
$=\frac{C_{h}}{2}\biggl{(}m_{k}(\tau_{k}\geq j+N_{\#}+1)+\ldots+m_{k}(\tau_{k+j}\geq N_{\#}+1)\biggr{)}.$
[…]
$(T_{k,k+j-1})_{*}m_{k}(\tau_{k+j}\geq N_{\#})\leq C\sum_{j\geq N_{\#}}j^{-p},$
where $C$ is a constant independent of $k,j$. Hence (A.1) follows by choosing $N_{\#}$ sufficiently large.

### v2

@@
 As applications, we obtain new estimates on the rate of memory loss for random ergodic compositions of Pomeau–Manneville type intermittent maps and intermittent maps with unbounded derivatives.
-Date: October 28, 2024
+Date: September 19, 2025
 A.K.
@@
 is supported by JSPS via the project LEADER.
-The authors thank Wael Bahsoun for useful discussions, as well as Loughborough University and Tokai University for hospitality over the course of this work.
+The authors thank Wael Bahsoun for useful discussions, Loughborough University and Tokai University for hospitality over the course of this work, and anonymous reviewers for many helpful comments and questions.
 1. Introduction
@@
 The statistical theory of nonstationary dynamical systems characterizes the long term behavior of $x_{n}$, for example through limit theorems for partial sums $\sum_{k=0}^{n-1}v_{k}(x_{k})$ generated by observations $v_{k}:X\to{\mathbb{R}}$ along the time-dependent trajectory.
-[…]
+A classical topic in dynamical systems theory is the decay of correlations, which describes how quickly initial distributions converge toward an invariant distribution.
+However, a nonstationary dynamical system may lack any natural invariant distribution.
+In this context, memory loss, a counterpart to correlation decay, was studied by Ott, Stenlund, and Young [22].
+Memory is said to be lost if the time-evolutions $\mu_{n},\mu_{n}^{\prime}$ of any two sufficiently regular initial distributions $\mu_{0},\mu_{0}^{\prime}$ on $X$ satisfy $\|\mu_{n}-\mu_{n}^{\prime}\|\to 0$ as $n\to\infty$ with respect to a suitable notion of distance $\|\cdot\|$.
+This property can be interpreted as all regular measures being attracted by the same moving target in a space of measures.
+The rate of memory loss has been the subject of several studies in the dynamical systems literature, including [19, 1, 5, 27, 13], and it plays an important role in further statistical analysis of nonstationary models, as evidenced by studies such as [28, 14, 9, 17, 7, 8, 10, 21], among others.
 Intermittent dynamical systems are chaotic systems that include laminar regions where trajectories can become trapped for long periods of time before transitioning to more chaotic behavior, as seen in the Pomeau–Manneville scenario [24].
 The rate of memory loss for such systems, whether stationary or nonstationary, is known to be polynomial [1, 29, 18, 16].
-[…]
+We now briefly describe the contents of our study.
+Building on the coupling approach as in Korepanov, Kosloff, and Melbourne [15], in a previous study [16] we analyzed the rate of memory loss in nonstationary nonuniformly expanding dynamical systems.
+We considered an abstract framework where the time-dependent trajectory $T_{1,n}:=T_{n}\circ\cdots\circ T_{1}$ makes frequent returns to a reference set $Y\subset X$, with first return dynamics that exhibit “good” distortion properties.
+Under a polynomially decaying upper bound on the tails of first return times, we derived memory loss estimates of the form
 $|(T_{1,n})_{*}\mu-(T_{1,n})_{*}\nu|=O(n^{-\gamma})$ (1.1)
-[…]
+for suitable probability measures $\mu$ and $\nu$, where $|\cdot|$ is the total variation distance and $(\cdot)_{*}$ is the pushforward.
+The rate coincides with the sharp rate established by Gouëzel [11] in the stationary case, meaning that it cannot be improved without additional information about the sequence $(T_{n})$.
+On the other hand, the estimates in [16] represent a worst-case scenario where all maps satisfy uniform tail bounds of the same polynomial order.
+In contrast, the present manuscript derives estimates on the rate of memory loss assuming non-uniform tail bounds that can vary with time.
+This flexibility enables us to address questions such as how the rate of convergence in (1.1) is influenced by subsequences $(T_{n_{k}})$ of “good” maps, i.e.
+maps for which the decay rate of stationary correlations is fast in comparison with other maps in the sequence.
+For an illustration of this phenomenom, see Theorem 1.1 below.
 We apply our abstract result (Theorem 2.6) to obtain new bounds on memory loss for random ergodic compositions of intermittent interval maps.
@@
 In our analysis we exploit the distortion bounds proved in [20].
-001/2121/21 / 21111001111
-11-1- 100111111-1- 1001111
-11-1- 100111111-1- 1001111
+01/21/211011
+$1-1011$
+$1-1011$
 Figure 1. Graphs of three intermittent maps. Left: Cui’s map in (3.4); Middle: Pikovsky map (3.7); Right: Grossmann–Horner map (3.14).
@@
 Let $\beta\in(0,1)$ and $a_{\beta}>2^{\beta}(\beta+2)$. Following [18], we define the cone
-${\mathcal{C}}_{*}(\beta)=\{f\in C((0,1])\cap L^{1}\,:\,$ $f\geq 0$, $f$ decreasing,
-[…]
+${\mathcal{C}}_{*}(\beta)=\{f\in C((0,1])\cap L^{1}\,:\,$ $\geq 0, ff$ decreasing,
+$\text{$x^{\beta+1}f$ increasing, $f(x)\leq a_{\beta}x^{-\beta}\int_{0}^{1}f(x)\,dx$}\}.$
 This cone contains densities of invariant absolutely continuous probability measures for maps (1.2) with parameters $\gamma\in[0,\beta]$.
@@
 By [16, Theorem 1.1], whenever $\mu$ and $\nu$ are measures with Hölder continuous densities,
-$|(T_{1,n})_{*}(\mu-\nu)|=O(n^{-\gamma_{*}}),\quad\text{as $n\to\infty$.}$ (1.3)
-[…]
+$|(T_{1,n})_{*}(\mu-\nu)|=O(n^{-1/\gamma_{*}}),\quad\text{as $n\to\infty$.}$ (1.3)
+In other words, memory loss occurs at a rate corresponding to the map with parameter $\gamma_{*}$, which is optimal for a general sequence $(T_{n})$ with $\sup_{n}\gamma_{n}\leq\gamma_{*}$.
+However, the following result shows that if the upper and lower asymptotic densities of “good” maps in the sequence $(T_{1,n})$ are sufficiently close, this rate can be improved:
 Theorem 1.1.
 Let $\gamma\in(0,1)$. There exists a constant $\kappa_{0}\in(0,1)$ such that the following holds. Assume that for some $a>0$, $\kappa\leq\kappa_{0}$, and $N\in{\mathbb{N}}$,
-#$\frac{\#\{1\leq k\leq n\>:\>\gamma_{k}\leq\gamma\}}{n}\in[a(1-\kappa),a(1+\kappa)]$ (1.4)
+$\frac{\#\{1\leq k\leq n\>:\>\gamma_{k}\leq\gamma\}}{n}\in[a(1-\kappa),a(1+\kappa)]$ (1.4)
 holds for all $n\geq N$.
@@
 See Corollary 3.3 for a precise statement.
+Remark 1.4.
+An interesting question, for which we do not have an answer, is whether the assumption on the upper asymptotic density of good maps is necessary in Theorem 1.1.
+It is important for our arguments, however: it makes sure that the good maps are distributed sufficiently uniformly (see the proof of Proposition 3.1), for which the lower bound alone is not sufficient.
 Structure of the paper
@@
 with a convention that $\log 0=$-$\infty$ and $\log 0-\log 0=0$.
-We suppose that there exist constants $\lambda>1$, $K>0$, $\delta_{0}>0$ and $n_{0}\geq 1$ such that the following assumptions hold.
+We suppose that there exist constants $\lambda>1$ and $K>0$ such that the following assumptions hold.
 For $x\in X$, $k\geq 1$, define return time functions $\tau_{k}:X\to\{1,2,\ldots\}$,
@@
   3. (NU:3)
-    If $a\subset Y_{k}$, then the map $F_{a}=T_{k,k+\tau_{k}(a)-1}:a\to Y_{k+\tau_{k}(a)}$ is a bijection, and for all $y,y^{\prime}\in a$,
+    If $a\subset Y_{k}$, then the map $F_{a}=T_{k,k+\tau_{k}(a)-1}|_{a}:a\to Y_{k+\tau_{k}(a)}$ is a bijection, and for all $y,y^{\prime}\in a$,
     $d(F_{a}(y),F_{a}(y^{\prime}))\geq\lambda d(y,y^{\prime}).$
@@
   4. (NU:4)
-    For all $x,x^{\prime}\in a$, with $F_{a}=T_{k,k+\tau_{k}(a)-1}$ as above,
+    For all $x,x^{\prime}\in a$, with $F_{a}=T_{k,k+\tau_{k}(a)-1}|_{a}$ as above,
     $\max_{0\leq j\leq\tau_{k}(a)}d(T_{k,k+j-1}(x),T_{k,k+j-1}(x^{\prime}))\leq Kd(F_{a}(x),F_{a}(x^{\prime})).$
@@
     there exist $\delta_{\#}>0$ and coprime integers $p_{1},\ldots p_{M}$ such that $m_{k}(\tau_{k}=p_{m})\geq\delta_{\#}$ for all $k\geq 1$ and $1\leq m\leq M$.
-[…]
+Moreover, for condition (i) it is sufficient that $\sup_{k\geq 1}\int\tau_{k}^{p}\,dm_{k}<\infty$ holds for some $p>1$.
 Proof.
-The proof is similar to in the stationary case [15, Section 4.1]. For completeness, we provide more details in Appendix A. ∎
+The proof is similar to that in the stationary case [15, Section 4.1]. For completeness, we provide more details in Appendix A. ∎
 2.2. Memory loss
@@
 3.1. Liverani–Saussol–Vaienti maps
-We consider sequences $(T_{n})$ of maps in the family (1.2).
+We consider sequences of maps $T_{n}:X\to X$ in the family (1.2) with $X=[0,1]$.
 First, we prove Theorem 1.1 concerning the rate of memory loss for deterministic compositions, after which we look at applications for random compositions.
@@
 Let $m$ be the Lebesgue measure on $Y=[1/2,1]$ normalized to probability. Denote by $g_{k}:X\to[0,1/2]$ and $h:X\to Y$ respectively the left and right inverse branches of $T_{k}$. Set
-$x_{n}(k)=g_{k}\cdots g_{k+n-1}(1)\quad\text{and}\quad y_{n}(k)=hg_{k+1}\cdots g_{k+n-1}(1).$
-Then, for all $k\geq 1$ and $n\geq 1$, $T_{k,k+n-1}$ maps both $[x_{n+1}(k),x_{n}(k)]$ and $[y_{n+1}(k),y_{n}(k)]$ bijectively onto $Y$. For each $k\geq 1$, we define a partition of $X$ by
+$x_{n}(k)=(g_{k}\circ\cdots\circ g_{k+n-1})(1)\quad\text{and}\quad y_{n}(k)=(h\circ g_{k+1}\circ\cdots\circ g_{k+n-1})(1).$
+Then, for all $k\geq 1$ and $n\geq 1$, $T_{k,k+n-1}$ maps both $x_{n+1}(k),x_{n}(k)]$ and $y_{n+1}(k),y_{n}(k)]$ bijectively onto $Y$. For each $k\geq 1$, we define a partition of $X$ by
 ${\mathcal{P}}_{k}=\{[x_{n+1}(k),x_{n}(k)]\>:\>n\geq 1\}\cup\{[y_{n+1}(k),y_{n}(k)]\>:\>n\geq 1\}.$
@@
 Let $\mu$ be a probability measure with density in ${\mathcal{C}}_{*}(\gamma)$.
-Then $\mu$ has tail bound $r(n)=O(n^{1-1/\gamma})$ with respect to $T_{1},T_{2},\ldots$ Moreover, for each $k\geq 1$, $h^{k}$ in (2.1) satisfies $h^{k}(\ell)\leq C(1\vee(\ell-\Theta k))^{-1/\gamma}$ where $C$ is a constant independent of $k$, and $\Theta=4\kappa(1-\kappa)^{-1}$.
+Then $\mu$ has tail bound $r(n)=O(n^{1-1/\gamma})$ with respect to $T_{1},T_{2},\ldots$ Moreover, for each $k\geq 1$, $h^{k}$ in (2.1) satisfies $h^{k}(\ell)\leq C(1\vee(\ell-\Theta k))^{-1/\gamma}$ where $C$ is a constant independent of $k$, and $\Theta=2\kappa(1-\kappa)^{-1}$.
 Proof.
-Let $\gamma_{n_{1}},\gamma_{n_{2}},\ldots$ be the sub-sequence of all parameters $\gamma_{n_{j}}$ with $\gamma_{n_{j}}\leq\gamma$, which by (3.2) must be an infinite sequence.
+Let $\gamma_{n_{1}},\gamma_{n_{2}},\ldots$ be the subsequence of all parameters $\gamma_{n}$ with $\gamma_{n}\leq\gamma$, which by (3.2) must be an infinite sequence.
 Let $\widetilde{T}_{j}=T_{n_{j}}$.
@@
 Next, for $k,\ell\geq 1$, denote
-$G_{\ell}(k)$ =#$=\#\{k\leq j\leq k+\ell-1\>:\>\gamma_{j}\leq\gamma\}=\#\bigl{(}\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr{)}.$
+$G_{\ell}(k)$ $=\#\{k\leq j\leq k+\ell-1\>:\>\gamma_{j}\leq\gamma\}=\#\bigl(\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr).$
 Recall that $\tau_{k}$ denotes the first return time to $Y$ w.r.t. $T_{k},T_{k+1},\ldots$ We observe that
@@
 $G_{\ell}(k)$ $=G_{k+\ell-1}(1)-G_{k-1}(1)\geq(k+\ell-1)a(1-\kappa)-(k-1)a(1+\kappa)$
-$\geq\frac{a(1-\kappa)}{2}\biggl{(}\ell-\frac{4\kappa}{1-\kappa}k\biggr{)},$ (3.3)
-provided that $\ell\geq 4(1+\kappa)/(1-\kappa)$. ∎
+$\geq a(1-\kappa)\biggl(\ell-\frac{2\kappa}{1-\kappa}k\biggr).$ (3.3)
 Remark 3.2.
@@
 $T(x)=\begin{cases}x(1+2^{\gamma}x^{\gamma}),&x\in[0,\tfrac{1}{2}),\\ 2^{\beta}(x-\tfrac{1}{2})^{\beta},&x\in[\tfrac{1}{2},1].\end{cases}$ (3.4)
-The map preserves an absolutely continuous probability measure if and only if $\gamma\beta<1$.
+The map $T$ preserves an absolutely continuous probability measure if and only if $\gamma\beta<1$.
 For sequences $T_{1},T_{2},\ldots$ of maps in (3.4) with parameters $\gamma_{k}\leq\gamma_{*}$ and $\gamma_{k}\leq\beta_{*}$ where $\gamma_{*}\beta_{*}<1$, an application of Theorem 2.6 yields the following estimate provided that (3.2) holds: for any two probability measures $\mu$ and $\mu^{\prime}$ with Hölder continuous densities,
@@
 3.1.1. Random compositions
-Let $\omega=(\omega_{n})$ be a random sequence of parameters sampled from a probability space $(\Omega,{\mathcal{F}},\mathbf{P})=(\Omega_{0}^{{\mathbb{N}}},{\mathcal{E}}^{{\mathbb{N}}},\mathbf{P})$, where $\Omega_{0}\subset(0,1)$ is an interval, ${\mathcal{E}}$ is the Borel sigma-algebra on $\Omega_{0}$, and =${\mathbb{N}}=\{1,2,\ldots\}$.
+Let $\omega=(\omega_{n})$ be a random sequence of parameters sampled from a probability space $(\Omega,{\mathcal{F}},\mathbf{P})=(\Omega_{0}^{{\mathbb{N}}},{\mathcal{E}}^{{\mathbb{N}}},\mathbf{P})$, where $\Omega_{0}\subset(0,1)$ is an interval, $\mathcal{E}$} is the Borel sigma-algebra on $\Omega_{0}$, and ${\mathbb{N}}=\{1,2,\ldots\}$.
 Denote by $T_{n}\circ\cdots\circ T_{1}$ compositions of maps $T_{k}$ along the random sequence $\omega$, where each $T_{k}$ corresponds to a map with parameter $\omega_{k}$.
@@
 The following result provides an example in the case of asymptotically mean stationary strongly mixing noise.
+Recall that, given two sub-$\sigma$-algebras $\mathcal{U},\mathcal{V}\subset\mathcal{{\mathcal{F}}}$, the $\alpha$-mixing coefficient between $\mathcal{U}$ and $\mathcal{V}$ is defined by
+$\alpha(\mathcal{U},\mathcal{V})=\sup_{U\in\mathcal{U},\>V\in\mathcal{V}}|\mathbf{P}(U\cap V)-\mathbf{P}(U)\mathbf{P}(V)|.$
 Proposition 3.7.
@@
     There exists a probability measure $\bar{\mathbf{P}}$ such that for any bounded measurable function $g:\Omega\to{\mathbb{R}}$,
-[…]
+    $\lim_{n\to\infty}\int_{\Omega}n^{-1}\sum_{k=0}^{n-1}g\circ\sigma^{k}\,d\mathbf{P}=\int_{\Omega}g\,d\bar{\mathbf{P}}.$
   • (B2)
@@
     $\alpha(n)=O(\log^{-\beta}(n))\quad\text{for some $\beta>1$.}$
-    Here, ${\mathcal{F}}_{1}^{i}$ is the sigma-algebra on $\Omega$ generated by the projections $\pi_{1},\ldots,\pi_{i}$; ${\mathcal{F}}_{i+n}^{\infty}$ is the sigma-algebra generated by $\pi_{i+n}\pi_{i+n+1},\ldots$, with $\pi_{k}(\omega)=\omega_{k}$; and
-    $\alpha({\mathcal{F}}_{1}^{i},{\mathcal{F}}_{j}^{\infty})=\sup_{A\in{\mathcal{F}}_{1}^{i},\,B\in{\mathcal{F}}_{j}^{\infty}}|\mathbf{P}(A\cap B)-\mathbf{P}(A)\mathbf{P}(B)|.$
+    Here, ${\mathcal{F}}_{1}^{i}$ is the sigma-algebra on $\Omega$ generated by the projections $\pi_{1},\ldots,\pi_{i}$; ${\mathcal{F}}_{i+n}^{\infty}$ is the sigma-algebra generated by $\pi_{i+n},\pi_{i+n+1},\ldots$, with $\pi_{k}(\omega)=\omega_{k}$.
   • (B3)
@@
 Proof.
-Define $\varphi$ as in the proof of Corollary 3.3. By (B3), it is enough to show that, for $\mathbf{P}$-a.e. $\omega\in\Omega$,
+Define $\varphi=\mathbf{1}_{[0,\gamma]}\circ\pi_{1}$ as in the proof of Corollary 3.3. By (B3), it is enough to show that, for $\mathbf{P}$-a.e. $\omega\in\Omega$,
 $\lim_{n\to\infty}n^{-1}\sum_{k=0}^{n-1}\varphi\circ\sigma^{k}=\bar{\mathbf{P}}(\omega_{1}\leq\gamma).$ (3.6)
 To this end, we decompose $n^{-1}\sum_{k=0}^{n-1}\varphi\circ\sigma^{k}-\bar{\mathbf{P}}(\omega_{1}\leq\gamma)=I+II$, where
-[…]
+$I=\int_{\Omega}n^{-1}\sum_{k=0}^{n-1}\varphi\circ\sigma^{k}\,d\mathbf{P}-\int_{\Omega}\varphi\,d\bar{\mathbf{P}}$
 and
-[…]
-By (B1), $I\to 0$ as $n\to\infty$.
-By (B2), $(\xi_{n})$ is a uniformly bounded sequence of $\alpha$-mixing random variables whose $\alpha$-mixing coefficients decay at the rate $O(\log^{-\beta}(n))$ with $\beta>1$.
+$II=n^{-1}\sum_{k=0}^{n-1}\xi_{k},\quad\xi_{k}=\varphi\circ\sigma^{k}-\int_{\Omega}\varphi\circ\sigma^{k}\,d\mathbf{P}.$
+By (B1), $I\to 0$ as $n\to\infty$. Moreover, $\sup_{k}\sup_{\omega\in\Omega}|\xi_{k}(\omega)|\leq 2$ and, by (B2),
+$\alpha(\sigma(\xi_{i}:0\leq i\leq j),\sigma(\xi_{i}:i\geq j+n))\leq\alpha(n)=O(\log^{-\beta}(n)).$
+Here, $\sigma(\xi_{i}:0\leq i\leq j)$ and $\sigma(\xi_{i}:i\geq j+n)$ denote the sub-$\sigma$-algebras generated by $\{\xi_{i}\}_{0\leq i\leq j}$ and $\{\xi_{i}\}_{i\geq j+n}$, respectively.
+Thus, $(\xi_{n})$ is a uniformly bounded sequence of $\alpha$-mixing random variables with coefficients decaying at the rate $O(\log^{-\beta}(n))$ with $\beta>1$.
 It follows from [26, Corollary 1] that $II\to 0$ as $n\to\infty$, for $\mathbf{P}$-a.e.
@@
 3.2. Pikovsky maps
-For $\gamma>1$, let $T:X\to X$ be a map on $X=[-1,1]$ defined implicitly by the relations
+For $\gamma>1$, let $T:X\to X$ be a map on X=[1,1]X=[-1,1] defined implicitly by the relations
 $x=\begin{cases}\frac{1}{2\gamma}(1+T(x))^{\gamma},&0\leq x\leq\frac{1}{2\gamma},\\ T(x)+\frac{1}{2\gamma}(1-T(x))^{\gamma},&\frac{1}{2\gamma}\leq x\leq 1,\end{cases}$ (3.7)
-and by letting $T(x)=-T(-x)$ for $x\in[-1,0]$.
+and by letting T(x)=T(x)T(x)=-T(-x) for $x\in[-1,0]$.
 The graph of $T$ is illustrated in Figure 1.
-The map has neutral fixed points at $x=1,-1$, while at $x=0$ its derivative becomes infinite.
+The map has neutral fixed points at x=1,1x=1,-1, while at $x=0$ its derivative becomes infinite.
 There is an interplay between the neutrality of the fixed points and the degree of the singularity, which is controlled by the parameter $\gamma$:
-$T^{\prime}(x)\approx C_{\gamma}x^{1/\gamma-1}-1\quad\text{as $x\searrow 0$},$
+$T^{\prime}(x)\approx C_{\gamma}x^{1/\gamma-1}\quad\text{as $x\searrow 0$},$
 $T^{\prime}(x)\approx 1+\frac{1}{2}(1-x)^{\gamma-1}\quad\text{as $x\nearrow 1$}.$
@@
 Before proving Theorem 3.8, we state a corollary for random compositions.
-Let $(\omega_{n})$ be a random sequence of parameters sampled from a probability space $(\Omega,{\mathcal{F}},\mathbf{P})=(\Omega_{0}^{{\mathbb{N}}},{\mathcal{E}}^{{\mathbb{N}}},\mathbf{P})$, where $\Omega_{0}=[\gamma_{-},\gamma_{+}]$ with $31<\gamma_{-}<\gamma_{+}<$ 3, and ${\mathcal{E}}$ is the Borel $\sigma$-algebra on $\Omega_{0}$.
+Let $(\omega_{n})$ be a random sequence of parameters sampled from a probability space $(\Omega,{\mathcal{F}},\mathbf{P})=(\Omega_{0}^{{\mathbb{N}}},{\mathcal{E}}^{{\mathbb{N}}},\mathbf{P})$, where $\Omega_{0}=[\gamma_{-},\gamma_{+}]$ with $1<\gamma_{-}<\gamma_{+}<3$, and $\mathcal{E}$} is the Borel $\sigma$-algebra on $\Omega_{0}$.
 As in Section 3.1.1, denote by $T_{n}\circ\cdots\circ T_{1}$ compositions of maps $T_{k}$ from (3.7) along the random sequence $\omega$, where each $T_{k}$ corresponds to a map with parameter $\omega_{k}$.
@@
 Further details are provided below.
-Let $(T_{k})$ be a sequence of maps in the family (3.7) with parameters $\gamma_{k}\in[\gamma_{-},\gamma_{+}]$, where $31<\gamma_{-}<\gamma_{+}<$ 3.
+Let $(T_{k})$ be a sequence of maps in the family (3.7) with parameters $\gamma_{k}\in[\gamma_{-},\gamma_{+}]$, where $1<\gamma_{-}<\gamma_{+}<3$.
 Let $X_{-}=(-0,1)$ and $X_{+}=(0,1)$.
@@
 Then $\{\Delta_{n}^{-}(k)\}_{n\geq 1}$ is a (mod $\lambda$) partition of $X_{-}$, $\{\Delta_{n}^{+}(k)\}_{n\geq 1}$ is a (mod $\lambda$) partition of $X_{+}$, and $T_{k}$ maps $\Delta_{n}^{\pm}(k)$ bijectively onto $\Delta_{n-1}^{\pm}(k+1)$.
-Moreover, $\Delta^{+}_{\ell}(k)=(x_{\ell}^{+}(k),x_{\ell+1}^{+}(k))$ and $\Delta^{-}_{\ell}(k)=(x_{\ell+1}^{-}(k),x_{\ell}^{-}(k))$, where $x_{\ell}^{+}(k)=g_{k,+}\cdots g_{k+\ell-1,+}(0)$.
+Moreover, $\Delta^{+}_{\ell}(k)=(x_{\ell}^{+}(k),x_{\ell+1}^{+}(k))$ and $\Delta^{-}_{\ell}(k)=(x_{\ell+1}^{-}(k),x_{\ell}^{-}(k))$, where $x_{\ell}^{+}(k)=(g_{k,+}\circ\cdots\circ g_{k+\ell-1,+})(0)$ and $x_{\ell}^{-}(k)=(g_{k,-}\circ\cdots\circ g_{k+\ell-1,-})(0)$.
 Next, for each $k,n\geq 1$ define
@@
 By inspecting the definition of $T$ in (3.7), we see that $|F_{a}(y)-F_{a}(y^{\prime})|\geq\Lambda|y-y^{\prime}|>0$ holds for all $y,y^{\prime}\in a\subset Y_{k}$ whenever $a\in{\mathcal{P}}_{k}$ where $F_{a}=T_{k,k+\tau_{k}(a)-1}$ and $\Lambda=\Lambda(\gamma_{-},\gamma_{+})>1$.
-Then (NU:3) follows by [20, Lemma 2.11] provided that $31<\gamma_{-}<\gamma_{+}<$ 3:
+Then (NU:3) follows by [20, Lemma 2.11] provided that $1<\gamma_{-}<\gamma_{+}<3$:
 Proposition 3.11.
-Assume that $31<\gamma_{-}<\gamma_{+}<$ 3. Then there exists $K>0$ such that for any $k\geq 1$ and $a\subset Y_{k}$, $a\in{\mathcal{P}}_{k}$,
+Assume that $1<\gamma_{-}<\gamma_{+}<3$. Then there exists $K>0$ such that for any $k\geq 1$ and $a\subset Y_{k}$, $a\in{\mathcal{P}}_{k}$,
 $\zeta=\frac{d(F_{a})_{*}(m_{k}|_{a})}{dm_{k+\tau_{k}(a)}}\quad\text{satisfies}\quad|\zeta|_{{\mathrm{LL}},k+\tau_{k}(a)}\leq K.$
@@
 In this case we can exploit the trick described in [4, Remark 1].
-Namely, by symmetry of $T_{k}$ we have $-z^{\prime}\in\delta_{n}^{-}(k)$ and $T_{j}^{\prime}(y)=T_{j}^{\prime}(-x)$.
+Namely, by symmetry of $T_{k}$ we have $z^{\prime}\in\delta_{n}^{-}(k)$ and $T_{j}^{\prime}(y)=T_{j}^{\prime}(-x)$.
 Since each $T_{j}$ is an odd function,
@@
 where the last inequality is due to [4].
-Hence, $\lambda(\cup_{\ell\geq n}\Delta_{\ell}^{+}(k))=\lambda((x_{n-1}(k)^{+},1))=O(n^{-1/(\gamma-1)})$.
+Hence, $\lambda(\cup_{\ell\geq n}\Delta_{\ell}^{+}(k))=\lambda((x_{n-1}^{+}(k),1))=O(n^{-1/(\gamma-1)})$.
 The same estimate holds with $\Delta_{\ell}^{-}(k)$ in place of $\Delta_{\ell}^{+}(k)$, leading to $\lambda(\cup_{\ell\geq n}\Delta_{\ell}(k))=O(n^{-1/(\gamma-1)})$.
@@
 Next, let $(T_{j})$ be a sequence of maps as in Theorem 3.8.
-Similar to the proof of Theorem 1.1, we let $\widetilde{T}_{j}=T_{n_{j}}$ where $\gamma_{n_{1}},\gamma_{n_{2}},\ldots$ is the subsequence of all parameters $\gamma_{n_{j}}$ with $\gamma_{n_{j}}\leq\gamma$ and, for each $k,\ell\geq 1$, define
-$G_{\ell}(k)$ =#$=\#\{k\leq j\leq k+\ell-1\>:\>\gamma_{j}\leq\gamma\}=\#\bigl{(}\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr{)}.$
+Similar to the proof of Theorem 1.1, we let $\widetilde{T}_{j}=T_{n_{j}}$ where $\gamma_{n_{1}},\gamma_{n_{2}},\ldots$ is the subsequence of all parameters $\gamma_{n}$ with $\gamma_{n}\leq\gamma$ and, for each $k,\ell\geq 1$, define
+$G_{\ell}(k)$ $=\#\{k\leq j\leq k+\ell-1\>:\>\gamma_{j}\leq\gamma\}=\#\bigl(\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr).$
 Further, let
 $\widetilde{\tau}_{k}(\ell)=\inf\{\ell\geq 1\>:\>\widetilde{T}_{m(k),m(k)+\ell-1}(x)\in Y_{n_{m(k)+\ell}}\},\quad m(k)=\min\{j\geq 1\>:\>n_{j}\geq k\},$
-where $\widetilde{T}_{m(k),m(k)+\ell-1}=\widetilde{T}_{m(k)}\circ\cdots\circ\widetilde{T}_{m(k)+\ell-1}$. By the definitions of $\Delta_{n}(k)$ and $\delta_{n}(k)$, we have
+where $\widetilde{T}_{m(k),m(k)+\ell-1}=\widetilde{T}_{m(k)+\ell-1}\circ\cdots\circ\widetilde{T}_{m(k)}$. By the definitions of $\Delta_{n}(k)$ and $\delta_{n}(k)$, we have
 $\bigcup_{j\geq\ell}\Delta_{j}(k)\subset\{\widetilde{\tau}_{k}\geq G_{\ell}(k)\}\quad\text{and}\quad\delta_{j}(k)\subset T_{k}^{-1}\Delta_{j-1}(k+1).$
@@
 $\leq\mu(\widetilde{\tau}_{k}\geq G_{\ell}(k))+(T_{k})_{*}\mu(\widetilde{\tau}_{k+1}\geq G_{\ell-1}(k+1))$
-[…]
+$\leq C\lambda(\widetilde{\tau}_{k}\geq G_{\ell}(k))+C\lambda(\widetilde{\tau}_{k+1}\geq G_{\ell-1}(k+1))$
 $\leq C^{\prime}G_{\ell}(k)^{-\frac{1}{\gamma-1}}+C^{\prime}G_{\ell-1}(k+1)^{-\frac{1}{\gamma-1}},$
@@
 Let $a\in{\mathcal{P}}_{k}$. Denote by $\rho$ the density of $\mu$. First, assume that $\rho$ is Lipschitz continuous with $\inf\rho=b>0$. Then, with $F_{a}$ as in (NU:3),
-$\rho_{a}(y):=\frac{d(F_{a})_{*}(\mu|_{a})}{dm_{k+\tau_{k}(a)}}(y)\propto\frac{\rho(y_{a})}{F_{a}^{\prime}(y_{a})},$
-where $y_{a}=F_{a}^{-1}(y)$. Hence, for $y,z\in Y_{k+\tau_{k}(a)}$, we have
+$\rho_{a}(y):=\frac{d(F_{a})_{*}(\mu|_{a})}{dm_{k+\tau_{k}(a)}}(y)=C_{k,a}\frac{\rho(y_{a})}{F_{a}^{\prime}(y_{a})},$
+where $y_{a}=F_{a}^{-1}(y)$ and $C_{k,a}$ is a scaling constant. Hence, for $y,z\in Y_{k+\tau_{k}(a)}$, we have
 $|\log\rho_{a}(y)-\log\rho_{a}(y^{\prime})|\leq I+II,$
@@
 Finally, if $\rho$ is Hölder continuous with exponent $\eta\in(0,1]$, then $\mu$ is regular with respect to the metric $d_{\eta}(x,y)=|x-y|^{\eta}$.
-Therefore, we can deal with this case by simply changing the metric from $d(x,y)=|x-y|$ to $d_{\eta}$ in the definitions of Section 2.1.
+Therefore, we can deal with this case by simply changing the metric from d(x,y)=|xy|d(x,y)=|x-y| to $d_{\eta}$ in the definitions of Section 2.1.
 ∎
@@
 3.3. Grossmann–Horner maps
-We follow the representation in [20].
-Let $T:X\to X$ be a map on $X=[-1,1]$ with graph as in Figure 1, having two surjective and symmetric branches on the subintervals $[-1,0]$ and $[0,1]$, such that $T|_{X\setminus\{0\}}\in C^{1}$ and $T|_{X\setminus\{0,-1,1\}}\in C^{2}$, defined locally by
+We follow the presentation in [20].
+Let $T:X\to X$ be a map on X=[1,1]X=[-1,1] with graph as in Figure 1, having two surjective and symmetric branches on the subintervals [1,0][-1,0] and $[0,1]$, such that $T|_{X\setminus\{0\}}\in C^{1}$ and $T|_{X\setminus\{0,-1,1\}}\in C^{2}$, defined locally by
 $T(x)=\begin{cases}1-b|x|^{\eta}+o(|x|^{\eta}),&\text{if $x\in U_{0}(\gamma,\eta)$},\\ -x+a|x-1|^{\gamma}+o(|x-1|^{\gamma}),&\text{if $x\in U_{1_{-}}(\gamma,\eta)$},\\ x+a|x+1|^{\gamma}+o(|x+1|^{\gamma}),&\text{if $x\in U_{-1_{+}}(\gamma,\eta)$}.\end{cases}$ (3.14)
-Here, $a,b>0$, $\eta>0$ and $\gamma>1$ are parameters, $U_{0}(\gamma,\eta)$ is a small neighborhood of 00, and $U_{1_{-}}(\gamma,\eta),U_{-1_{+}}(\gamma,\eta)$ are small neighborhoods of $1,-1$, respectively.
+Here, $a,b>0$, $\eta>0$ and $\gamma>1$ are parameters, $U_{0}(\gamma,\eta)$ is a small neighborhood of 0, and $U_{1_{-}}(\gamma,\eta),U_{-1_{+}}(\gamma,\eta)$ are small neighborhoods of 1,11,-1, respectively.
 For $i=1,2$, the derivatives of $T$ are assumed to satisfy
@@
   • $|T^{\prime}(x)|>1$ for all $x\neq-1,1$,
-  • $T$ is strictly monotonous and convex on the two intervals $[-1,0]$ and $[0,1]$, and
+  • $T$ is strictly monotonous and convex on the two intervals [1,0][-1,0] and $[0,1]$, and
   • the error terms in (3.14) and (3.3) are uniformly bounded.
@@
 The memory loss bound below follows by applying Theorem 2.6 in combination with results in [20].
-Fix $0<a_{-}<a_{+}<\infty$, $0<b_{-}<b_{+}<\infty$, $1<\gamma_{-}<\gamma_{+}\leq 2$, $10<\eta_{-}<\eta_{+}<$ 1.
+Fix $0<a_{-}<a_{+}<\infty$, $0<b_{-}<b_{+}<\infty$, $1<\gamma_{-}<\gamma_{+}\leq 2$, $0<\eta_{-}<\eta_{+}<1$.
 Let $T_{1},T_{2},\ldots$ be a sequence of maps in (3.14), such that the parameters $a_{i},b_{i},\eta_{i},\gamma_{i}$ associated to $T_{i}$ satisfy
@@
 3.3.1. Proof of Theorem 3.14
-Following the representation in [20], we start by describing several partitions associated with nonstationary compositions $T_{n}\circ\cdots\circ T_{1}$ of maps in (3.14).
-Let $X=[-1,1]$ and, as in Section 3.2.1, denote by $g_{k,-}:(-1,1)\to X_{-}$ and $g_{k,+}:(-1,1)\to X_{+}$ the left and right inverse branches of $T_{k}$, respectively.
+Following the presentation in [20], we start by describing several partitions associated with nonstationary compositions $T_{n}\circ\cdots\circ T_{1}$ of maps in (3.14).
+Let X=[1,1]X=[-1,1] and, as in Section 3.2.1, denote by $g_{k,-}:(-1,1)\to X_{-}$ and $g_{k,+}:(-1,1)\to X_{+}$ the left and right inverse branches of $T_{k}$, respectively.
 Here $X_{-}=(-1,0)$ and $X_{+}=(0,1)$.
@@
 $m_{k}(\tau_{k}=2)\geq|\delta_{2}^{-}(k)|\geq\delta_{\#}\quad\text{and}\quad m_{k}(\tau_{k}=3)\geq|\delta_{3}^{-}(k)|\geq\delta_{\#}.$
-[…]
+Moreover, $\sup_{k}\int\tau_{k}^{p}\,dm_{k}\leq\sup_{k}\lambda(Y_{k})^{-1}\int\tau_{k}^{p}\,d\lambda<\infty$ holds for some $p>1$ by Proposition 3.18 below, since $\eta_{+}(\gamma_{+}-1)<1$.
+Now, (NU:5) follows by Proposition 2.1.
 Proposition 3.18.
@@
 Set
-$G_{\ell}(k)$ =#($=\#\bigl{(}\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr{)},$
+$G_{\ell}(k)$ $=\#\bigl(\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr),$
 $\widetilde{\tau}_{k}(\ell)$ $=\inf\{\ell\geq 1\>:\>\widetilde{T}_{m(k),m(k)+\ell-1}(x)\in Y_{n_{m(k)+\ell}}\},$
@@
 We prove (a) first. Fix $n\geq 0$ and $k\geq 1$. For each $0\leq j\leq n$, define
-$Y^{\prime}_{j}$ =$=\{y\in Y_{k+n-j}:T_{k+n-j,k+\ell-1}(y)\notin Y_{k+\ell}\text{ for all $n-j<\ell\leq n$}\},$
+$Y^{\prime}_{j}$ $=\{y\in Y_{k+n-j}:T_{k+n-j,k+\ell-1}(y)\notin Y_{k+\ell}\text{ for all $n-j<\ell\leq n$}\},$
 $Y_{j}^{\prime\prime}$ $=Y_{k}\cap T_{k,k+n-j-1}^{-1}(Y_{j}^{\prime})$
-=$=\{y\in Y_{k}:T_{k,k+n-j-1}(y)\in Y_{k+n-j}\text{ and $T_{k,k+\ell-1}(y)\notin Y_{k+\ell}$ for all $n-j<\ell\leq n$}\}.$
+$=\{y\in Y_{k}:T_{k,k+n-j-1}(y)\in Y_{k+n-j}\text{ and $T_{k,k+\ell-1}(y)\notin Y_{k+\ell}$ for all $n-j<\ell\leq n$}\}.$
 In particular, $Y_{0}^{\prime}=Y_{k+n}$ and $Y^{\prime\prime}_{0}=\{y\in Y_{k}:T_{k,k+n-1}(y)\in Y_{k+n}\}$.
@@
 $\mu=\theta\mu_{n}+(1-\theta)\mu^{\prime}_{n},$
-where $\mu_{n}$, $\mu^{\prime}_{n}$ are probability measures with $(T_{k,k+n-1})_{*}\mu_{n}=m_{k+n}$ and is $(T_{k,k+n-1})_{*}\mu^{\prime}_{n}$ regular with tail bound $h_{n-N}^{k+N}$ w.r.t.
+where $\mu_{n}$, $\mu^{\prime}_{n}$ are probability measures with $(T_{k,k+n-1})_{*}\mu_{n}=m_{k+n}$ and $(T_{k,k+n-1})_{*}\mu^{\prime}_{n}$ is regular with tail bound $h_{n-N}^{k+N}$ w.r.t.
 $T_{k+n},T_{k+n+1},\ldots$
@@
 By [15, Proposition 4.7] there exist nonnegative constants $\xi_{\ell,n}$, $1\leq n\leq\ell<\infty$, such that
-[…]
-[…]
+$\sum_{n=1}^{\ell}\xi_{\ell,n}\mu(a_{k,n})$ $=r(\ell)-r(\ell+1)$ for each $\text${ for each $\ell$,
+$\sum_{\ell=n}^{\infty}\xi_{\ell,n}$ $=1$ $\text{ for each }n.$
 For $j\geq n_{0}+1$, define $\chi_{j}=\sum_{n=1}^{j-n_{0}}\xi_{j-n_{0},n}\nu_{n}$. Then $\mu=\sum_{j=n_{0}+1}^{\infty}\chi_{j}$ and $\chi_{j}(X)=\alpha_{j}$. From (4.3), we have
@@
 ${\hat{r}}(n)=\min\{1,r(1),\ldots,r(n)\}.$
-This way, ${\hat{r}}$ is nonincreasing and ${\hat{r}}(1)=1$; for a probability measure, the tail bounds $r$ and ${\hat{r}}$ are equivalent. Similarly define ${\hat{h}}^{k}_{n}$.
+This way, $^{\hat{r}}$ is nonincreasing and ${\hat{r}}(1)=1$; for a probability measure, the tail bounds $r$ and $^{\hat{r}}$ are equivalent. Similarly define ${\hat{h}}^{k}_{n}$.
 Let $X_{1},X_{2},\ldots$ be random variables with values in $\{n_{0},n_{0}+1,\ldots\}$, such that for all $\ell\geq n_{0}$,
@@
 To this end, fix $b\in(1/2,1)$ by $b=\bigl(1+(1-\theta)^{\frac{1}{2\beta^{\prime}}}\bigr)/2$ and let $A\geq 1$ be a large integer whose value will be specified later.
-Restrict to $\Theta$ sufficiently small so that $01-b-\Theta b>$ 0.
+Restrict to $\Theta$ sufficiently small so that $1-b-\Theta b>0$.
 Let $C$ denote various constants which depend only on $A,n_{0},\theta,\Theta,C_{h},C_{\beta},\beta,\beta^{\prime}$.
@@
 $\operatorname{\mathbb{P}}(S\geq n)$ $=\sum_{j\geq 1}\operatorname{\mathbb{P}}(S_{j}\geq n)\operatorname{\mathbb{P}}(\tau=j)$
-[…]
+$\leq C_{\beta}^{\prime}C_{2}(\Theta_{k}^{*}k+1)^{\beta^{\prime}}\theta(1-\theta)^{-1}\sum_{j\geq 1}(1-\theta)^{j/2}n^{-\beta^{\prime}}\leq C_{\beta}^{\prime}C_{3}(\Theta_{k}^{*}k+1)^{\beta^{\prime}}n^{-\beta^{\prime}},$
 as desired. ∎
@@
 $(T_{k,k+n-1})_{*}m_{k}(Y_{k+n})$
-[…]
+$\geq\sum_{\ell\leq N_{\#}}(T_{k+n-N,k+n-1})_{*}(\nu|_{\{\tau_{k+n-N}=\ell\}})(Y_{k+n})$
 $=\sum_{\ell\leq N_{\#}}(T_{k+n-N+\ell,k+n-1})_{*}(T_{k+n-N,k+n-N+\ell-1})_{*}(\nu|_{\{\tau_{k+n-N}=\ell\}})(Y_{k+n})$
 $\geq\sum_{\ell\leq N_{\#}}e^{-\text{diam}(X)K_{1}}(T_{k+n-N,k+n-N+\ell-1})_{*}(\nu|_{\{\tau_{k+n-N}=\ell\}})(Y_{k+n-N+\ell})$
-[…]
+$\times(T_{k+n-N+\ell,k+n-1})_{*}m_{k+n-N+\ell}(Y_{k+n}).$
 As in [15, Proposition 4.4], using our assumption $m_{k}(\tau_{k}=p_{m})\geq\delta_{\#}$ together with Proposition 2.5, we find that for any $\ell\leq N_{\#}$,
@@
 whenever $n>N_{1}+N_{\#}$. Thus, condition (NU:5) holds.
-[…]
+Finally, to see that $\sup_{k}\int\tau_{k}^{p}\,dm_{k}<\infty$ implies (A.1), suppose that $j,k\geq 1$.
+Since $m_{k}$ is a measure on $Y_{k}$ with $|m_{k}|_{{\mathrm{LL}},k}\leq K_{2}$, arguing as in the proof of Proposition 4.1(a), write
 $(T_{k,k+j-1})_{*}m_{k}(\tau_{k+j}>N_{\#})$ $\leq\frac{1}{2}h_{j}^{k}(N_{\#}+1)$
 $=\frac{C_{h}}{2}\biggl(m_{k}(\tau_{k}\geq j+N_{\#}+1)+\ldots+m_{k}(\tau_{k+j}\geq N_{\#}+1)\biggr).$
-[…]
+Using $\sup_{k}\int\tau_{k}^{p}\,dm_{k}<\infty$ together with Markov’s inequality, we obtain
 $(T_{k,k+j-1})_{*}m_{k}(\tau_{k+j}\geq N_{\#})\leq C\sum_{j\geq N_{\#}}j^{-p},$